### v1

# MACTAS: Self-Attention-Based Module for Inter-Agent Communication in Multi-Agent Reinforcement Learning

Abstract
Communication is essential for the collective execution of complex tasks by human agents, motivating interest in communication mechanisms for multi-agent reinforcement learning (MARL). However, existing communication protocols in MARL are often complex and non-differentiable. In this work, we introduce a self-attention-based communication module that exchanges information between the agents in MARL. Our proposed approach is fully differentiable, allowing agents to learn to generate messages in a reward-driven manner. The module can be seamlessly integrated with any action-value function decomposition method and can be viewed as an extension of such decompositions. Notably, it includes a fixed number of trainable parameters, independent of the number of agents. Experimental results on the SMAC benchmark demonstrate the effectiveness of our approach, which achieves state-of-the-art performance on several maps.
1 Introduction
Multi-agent reinforcement learning (MARL) (Ning and Xie 2024; Du and Ding 2021; Nguyen, Nguyen, and Nahavandi 2020; Oroojlooy and Hajinezhad 2023; Wong et al. 2023) promises efficient control in large-scale systems through the collective learning of many agents in dynamic environments. Multiple agents interact within a shared environment to learn optimal policies. MARL is defined as the process in which agents optimize their behaviors based on the history of their interactions with the environment, i.e., on their experience. Research in this area aims to design methods that enable agents to learn quickly, i.e., from minimal experience. The challenges and complexities in using MARL for controlling multi-agent systems include coordination among agents, non-stationary environments due to evolving agent policies, and scalability issues. Advanced MARL approaches often involve sophisticated techniques such as decentralized training, communication protocols among agents, and hierarchical methods to manage multi-agent interactions as effectively as possible. Theoretical advancements and practical applications of MARL demonstrate its potential in diverse fields ranging from robotics (Orr and Dutta 2023) and autonomous driving (Yadav, Mishra, and Kim 2023; Schmidt et al. 2022; Althamary, Huang, and Lin 2019) to telecommunications (Li et al. 2022b; Feriani and Hossain 2021).
Figure 1: 3s5z_vs_3s6z scenario in the StarCraft Multi-Agent Challenge (SMAC).
The most prominent research topics in MARL are the action-value function decompositions (Wang et al. 2021; Liu, Zhu, and Chen 2023), where agents share a joint reward according to their contribution, and inter-agent communication (Liu et al. 2023; Guo, Shi, and Fan 2023), where agents exchange information in order to coordinate their joint policy. The most challenging multi-agent environments are partially observable, usually via limiting agents’ information to their close proximity, and the successful achievement of their common goals is predicated upon coordination, which needs to involve some form of communication. In most research, the action-value function and inter-agent communication are issues addressed separately. Therefore, in this paper, we introduce communication into the very core of the action-value function decomposition.
The prevailing paradigm in cooperative MARL is Centralized Training and Decentralized Execution (CTDE), where the training algorithm utilizes the global state, but agents can only use their own local observations. We argue that in most challenging MARL tasks, this is insufficient. Moreover, in most challenging MARL tasks, some form of inter-agent communication is a necessity, which can be provided relatively easily (e.g., through WiFi). Therefore, we introduce a communication protocol based on collecting information from each agent in a central entity and broadcasting the combined information to all agents. Because information aggregation is based on attention, our proposed architecture requires a constant number of weights plus an embedding for each agent. It is easy to apply using readily available and efficient implementations of the Transformer encoder layer.
The contribution of this paper may be summarized as follows:
  • We introduce Multi-Agent Communication via Transformer on Agents’ States (MACTAS) module for inter-agent communication in MARL. This module allows the agent to share their state of consciousness.
  • We verify MACTAS experimentally on several challenging maps from SMAC, giving excellent results and establishing state-of-the-art in some cases.
2 Related Work
Single-Agent Reinforcement Learning (SARL) (Sutton and Barto 2018) addresses trial-and-error learning of sequential decision making under uncertainty. In Multi-Agent Reinforcement Learning (MARL) (Ning and Xie 2024), a group of agents learns to act in the same dynamic environment, collectively modifying its state. There are various settings for MARL, depending on whether the agents operate synchronously or asynchronously, whether they have common or adverse goals, and other circumstances. In this paper, we focus on the synchronous and cooperative MARL, with a single reward value for all agents at every time instance.
MARL has been addressed by extending well-established SARL algorithms: Multi-Agent SAC (Wei et al. 2018), Multi-Agent PPO (Yu et al. 2022), Multi-Agent DDPG (Lowe et al. 2017).
Generally, cooperative MARL adopts a centralized training with a decentralized execution (CTDE) paradigm, which suffers from the global action-value function, whose complexity grows exponentially with the number of agents (Kraemer and Banerjee 2016). Therefore, action-value function decomposition is a fundamental problem in MARL. Existing decomposed architectures include modules assigned to agents. These modules are fed with agents’ observations and their previous actions and produce agents’ contributions to the architecture’s output. The best action for the agent is the one that produces the highest contribution. In VDN (Sunehag et al. 2018), these contributions are just added. In QMIX (Rashid et al. 2018), a neural module, the mixing network, produces its weighted sum with positive weights. In Weighted QMIX (Rashid et al. 2020), the architecture is trained with a special weight assigned to the collective actions deemed the best. In QTRAN (Son et al. 2019), an additional module fed with all the actions of all agents corrects the weighted sum of the agents’ contributions. In Qatten (Yang et al. 2020), the mixing network is based on attention. In CDS (Li et al. 2021), a component of the mixing network optimizes the amount of exploration undertaken by the agents. In DVD (Li et al. 2022a), the internal state of the modules assigned to agents is also passed to the mixing networks. In SHAQ (Wang et al. 2022), the mixing network is enriched by several components based on Shapley value theory. In QPLEX (Wang et al. 2021), a carefully engineered duplex dueling of pairs of contributions of agents is introduced into the mixing network. In NA2Q (Liu, Zhu, and Chen 2023), a similar duplex dueling is performed with the use of attention.
Another fundamental problem in the cooperative MARL is the coordination of agents’ activity with a certain communication protocol. A number of solutions have been proposed for that purpose, including CommNet (Sukhbaatar, Szlam, and Fergus 2016) which introduces differentiable communication by aggregating a learned broadcast vector from each agent, TarMAC (Das et al. 2019) uses signature-based soft attention to direct agent communication towards recipients, NDQ (Wang et al. 2019) applies two forms of regularization – mutual information maximization and message entropy minimization to optimize message content, GA-Comm (Liu et al. 2020) uses two-stage attention mechanism to limit number of messages exchanged between agents, IS (Kim, Park, and Sung 2021) encodes trajectory of future actions for each agent and shares them with others, in MAIC (Yuan et al. 2022) agents explicitly model their teammates and use it to bias their value function, DHCG (Liu et al. 2023) learns a dependency graph between agents and uses it to limit the number of communication edges, TEM (Guo, Shi, and Fan 2023) which uses a Transformer to simulate communication resembling email exchange, where agents only directly communicate with teammates nearby, but can forward messages from one to another, and ACE (Li et al. 2023) which proposes bidirectional action-dependent Q-learning.
In addition, there are solutions in which the agents communicate in natural language, including: TWOSOME (Tan et al. 2024) which uses an LLM to evaluate the probabilities of joint action and applies LoRa for efficient PPO training, and Verco (Li et al. 2024) which separates the output of the LLM into communication and actions, allowing LLM agents to exchange interpretable verbal messages.
In our setting, we consider communication methods that can be connected with an arbitrary action-value decomposition method. We also allow agents to send messages based only on their local information (current observation, historical observations, and historical actions). In this framework, the best results so far have been achieved by MAIC (Yuan et al. 2022).
3 Method
Figure 2: The proposed MACTAS architecture.
3.1 Problem definition
In this paper, we discuss the problem of MARL using the formalism of a cooperative Decentralized Partially Observable Markov Decision Process (DecPOMDP) with communication. A DecPOMDP is defined by a tuple, $\langle\mathbb{N},\mathbb{S},\mathbb{A},\mathcal{P},\mathcal{O},\mathcal{R}\rangle$, where $\mathbb{N}=\{1,\dots,n\}$ is a team of agents, $\mathbb{S}$ is a space of environmental states, $\mathbb{A}=\mathbb{A}^{1}\times\dots\times\mathbb{A}^{n}$ is a space of team actions of the agents, $\mathcal{P}$ is the state transition probability, $\mathcal{O}$ is the observation function, and $\mathcal{R}$ is the reward function. At time $t=1,2,\dots$ each $i$-th agent makes an observation, $o^{i}_{t}=\mathcal{O}^{i}(s_{t})$, of the environment state $s_{t}$. Agents may exchange messages with others through a certain channel that may generally refine these messages with some processing.
Each agent maintains a representation, $\mathcal{I}^{i}_{t}$, of observations it has made, messages it has received, and actions it has taken. Based on thereof and a policy, $\pi^{i}$, the agent performs an action, $a^{i}_{t}\in\mathbb{A}^{i}$, $a^{i}_{t}\sim\pi^{i}(\cdot|\mathcal{I}^{i}_{t})$. The team action $a_{t}=[(a^{1}_{t}),\dots,(a^{n}_{t})]$ impacts the next environment state $s_{t+1}\sim\mathcal{P}(\cdot|s_{t},a_{t})$ and the reward $r_{t}=\mathcal{R}(s_{t+1})$ is given collectively to all agents.
The goal of the agents’ learning is to optimize their team policy $\pi=[\pi^{1},\dots,\pi^{n}]$ and the message exchange protocol for the team to expect in each environment state $s$ the highest sum of discounted rewards
$\mathcal{V}^{\pi}(s)=E\left(\sum_{i\geq 0}\gamma^{i}r_{t+i}\Big{|}s_{t}=s,\text{policy in use}=\pi\right),$
where $\gamma\in[0,1]$ is the discount factor.
The goal of this paper is to establish a message exchange protocol that supports the above goal of learning.
3.2 MACTAS
Our proposed Multi-Agent Communication via Transformer on Agents States (MACTAS) architecture is presented in Fig. 2. It is built on a decomposition of the action-value function. Each such decomposition includes the input modules assigned to the agents ($f^{i}$ assigned to the $i$-th agent for $i=1,\dots,n$ in Figure 2). These modules are fed with previous actions of these agents, $a^{i}_{t-1}$, and their current observations, $o^{i}_{t}$. They produce agents’ contributions, $\widehat{Q}^{i}_{t}$, to future discounted rewards and perhaps some other vectors (Li et al. 2022a). The outputs of the $f^{i}$ modules are fed to the mixing network that produces the value of the action-value function, $\widehat{Q}_{t}$. The input module $f^{i}$ usually comprises a feedforward part, denoted by MLP in Figure 2, a recurrent part, denoted by GRU, and the output feedforward part, again denoted by MLP.
Intuitively, we understand the $f^{i}$ module as the consciousness of the $i$-th agent. Based on what the agent is aware of, it determines its contribution to future discounted rewards, depending on the action it currently takes. In order to give a global view on the current state to the agent’s consciousness, we introduce a communication mechanism. We divide the $f^{i}$ module after its recurrent part and feed its state $h^{i}_{t}\in\mathbb{R}^{n_{h}}$ to the module $g$ where the states of consciousness of the agents are confronted. As a result, the $g$ module produces increments, $z^{i}_{t}\in\mathbb{R}^{n_{h}}$. The sum $z^{i}_{t}+h^{i}_{t}$ is then passed to the second MLP part of the $f^{i}$ module.
Consequently, the consciousness of each agent is based not only on its observations and previous actions, but also on observations and previous actions of all other agents.
The minimal requirement for the $g$ module is as follows:
  1. 1.
    It is fed with $n$ vectors and produces $n$ vectors, both of size $\dim(h^{i}_{t})$. $i$-th input vector corresponds to the $i$-th output vector.
  2. 2.
    It is a layered structure and the layers include self-attention (Vaswani et al. 2017).
Introducing $g$ satisfying the above requirements in their basic form is enough to outperform MARL without communication by a large margin. However, what follows from our experiments is that usually the best performing form of the $g$ module is a stack of Transformer (encoder) modules (Vaswani et al. 2017), with the following settings:
  1. 1.
    There are no positional embeddings. The agents’ embeddings are fed to $f^{i}$ instead.
  2. 2.
    There is no masking.
  3. 3.
    We use a separate Adam optimizer (Kingma and Ba 2017) for the Transformer parameters; the rest of the architecture is trained with its originally proposed optimizer (typically it is RMSProp: Tieleman and Hinton 2012).
  4. 4.
    The weights of the output layer of the top Transformer module are initialized with zeros. Therefore, initially, this module simply passes the input forward.
Good performance of the Transformer in this role is an expected result. This architecture has been specifically optimized to transform a set of embeddings into another set of the same size, which is exactly what we require in the $g$ module.
3.3 Implementation considerations
The basic approach to implement a multi-agent system with our proposed MACTAS module is to set up the Intermediate Processing Unit (IPU) on which the module is run. At each time instant, all $n$ agents transmit their internal states $h^{i}_{t}$ to the IPU. Then, the Transformer layers process all the input vectors and transmit the increments $z^{i}_{t}$ back to the agents. The number of operations necessary to process the Transformer modules is $O(n^{2})$. However, they are very efficiently parallelizable. The total size of all the transmissions is $O(n)$, specifically, equal to $2n\dim(h^{i}_{t})$.
Usually in multi-agent systems, reducing the communication capacity to $O(n)$ and offloading the communication processing burden to the IPU is very convenient. In these systems, the agents are naturally connected through a network and their computational capacity is limited (e.g., they are mobile and battery powered), contrary to that of the network infrastructure. Still, these are the agents that need to take on the processing of their observations and control.
In most inter-agent communication protocols, including MAIC (Yuan et al. 2022), all pairs of agents need to be able to communicate, which translates into $O(n^{2})$ messages. Technically, the parts of the agents that communicate with others could be moved to an IPU. That would require $O(n)$ messages to and from the IPU. Still, natural parallelization of MACTAS is its noticeable advantage.
A distributed implementation of a multi-agent system with our proposed module is also possible. Then, each agent needs to take on the part of processing that consists of refining $h^{i}_{t}$ and results in $z^{i}_{t}$. That requires exchange of messages with other agents as many times as many Transformer layers are stacked in MACTAS, each time with $O(n)$ messages; $O(n^{2})$ for all agents. In training, the agents need to aggregate increments in their shared weights.
In the distributed implementation, it may be the case that some of the agents are out of reach of others and do not exchange information with them. That only means that self-attention in MACTAS is forced to assign zero weight to unavailable transmissions from those unheard agents.
3.4 Exploration
Learning in a multi-agent system with decentralized execution is hampered by agents’ policies falling into suboptimal Nash equilibria. Let us consider a typical MARL setting where agents use the $\epsilon$-greedy strategy with $\epsilon$ annealing and explore independently. On average, only $n\epsilon$ agents explore. In order to learn that a certain combination of agents’ actions is better than the current Nash equilibrium, at least two agents need to explore and try a better combination. However, this is unlikely. What is more likely is that only one agent explores, which leads to a combination worse than greedy. Note that this effect may be even stronger for a better approximation of the $Q$ function, as it leaves less room for accidental switches of the greedy action combinations.
To counteract this effect, we propose a combination of $\epsilon$-greedy and Boltzmann exploration. Namely, each agent with probability $\epsilon$ chooses independently a random action; however, with probability $1-\epsilon$ it does not act greedily, but chooses from the top $k$ of its actions (according to its local action-value function) using the Boltzmann strategy. Using $k=1$ is equivalent to the pure $\epsilon$-greedy policy. Selecting a small $k$ leads to exploring the region that is quite good according to the agent. We observe that $k=2$ works well in practice. The temperature of the Boltzmann exploration controls the amount of extra exploration that we want to add. We plan to extend this idea and other exploration strategies for MARL in future work.
Hyperparameter Value
Batch size 32
Test episodes 32
Replay buffer size 5000
Discount factor 0.99
Start epsilon 1.0
Finish epsilon 0.05
Anneal steps 50,0001 or 100,0002
RNN units 64
Optimizer RMSProp
Learning rate 0.0005
Table 1: Basic experimental settings and hyperparameters. 1 for 2.5 mln. step maps. 2 for 7.5 mln. step maps.
Scenario Stacked encoders FFN dimension
2c_vs_64zg 2 512
5m_vs_6m 1 128
27m_vs_30m 1 128
MMM2 3 256
6h_vs_8z 4 512
3s5z_vs_3s6z 3 512
Table 2: MACTAS shape for each scenario.
Figure 3: Percentage of wins in test games in training time for the connection protocols MACTAS and MAIC, and the $Q$ architectures QMIX, QPLEX, and VDN.
Figure 4: The impact of our exploration scheme on MACTAS+QMIX.
Figure 5: The impact of our residual connection on MACTAS+QMIX.
4 Experiments
In this section, we present the results of our proposed MACTAS communication protocol in the StarCraft Multi-Agent Challenge (SMAC: Samvelyan et al. 2019) environment. We compare it with the state-of-the-art MAIC (Yuan et al. 2022), in combinations with 3 mixer algorithms: QMIX, QPLEX, VDN, and either MAIC or MACTAS communication protocol, on 6 maps. We present results of two ablation studies regarding the design of neural network connectivity and the exploration form. Additionally, in the Appendix, we report the sensitivity study of MACTAS to hyperparameters, lower-level SMAC settings, and analyses of the most interesting strategies that have emerged in training.
4.1 StarCraft Multi-Agent Challenge
StarCraft Multi-Agent Challenge (SMAC) is a micromanagement benchmark where two opposite teams face each other in different scenarios (maps), one team being controlled by the RL agent and the other by a built-in algorithm. Maps differ by difficulty, mainly due to an imbalance in the number of controlled units. For our experiments, we selected 2 hard and 4 super-hard scenarios, which are widely benchmarked in the MARL community:
  • 2c_vs_64zg with 2 colossus units leveraging terrain advantage against 64 small enemy zerglings. Level: hard.
  • 5m_vs_6m with 5 cooperating marines and 6 enemy marines. Level: hard.
  • 27m_vs_30m big battle of 27 marines conquering 30 marines. Level: super-hard.
  • MMM2 scenario with a diverse units containing 1 medivac (medic), 2 marauders, and 7 marines fighting team with 1 more marauder and 1 more marine. Level: super-hard.
  • 6h_vs_8z with 6 shooting hydralisks against 8 malee zealots. Level: super-hard.
  • 3s5z_vs_3s6z scenario with 3 stalkers and 5 zealots against 3 stalkers and 6 zealots. Level: super-hard.
In order to verify versatility of MACTAS algorithm, we run experiments on a diverse set of hard and super-hard scenarios. Every map provides the opportunity to find a unique winning strategy. The agent requires learning basic skills such as navigating the battlefield, managing health and shield, or focusing the fire. All considered scenarios are imbalanced in favor of the enemies; therefore, simple strategies are not sufficient to win the episode. Popular observed winning policies include crossing the obstacle on 2c_vs_64zg, flanking on 27m_vs_30m, ’kiting’ on 6h_vs_8z, ’hit and run’ on 3s5z_vs_3s6z (Samvelyan et al. 2019).
4.2 Experimental setting
We borrow the basic settings from NA2Q (Liu, Zhu, and Chen 2023) and extend their code with the implementation of MACTAS and MAIC. The algorithms are trained using the Double Q-Learning (Hasselt, Guez, and Silver 2016) algorithm. Each experiment is run for a fixed number of training episode steps – 2.5 million for 2c_vs_64zg, 5m_vs_6m, MMM2, 27m_vs_30m environments and 7.5 million for 6h_vs_8z, 3s5z_vs_3s6z environments, interleaved with test episodes every 2000 steps. Each experiment is repeated for seeds $1,\dots,5$. We report the average percentage of wins (and standard deviations in the Appendix) in test games.
All hyperparameters not strictly related to the communication algorithm are shared between maps, algorithms, and mixers and are presented in Table 1.
For MAIC communication, we use the same hyperparameters as in the original paper.
MACTAS hyperparameters
We notice that MACTAS is sensitive to communication hyperparameters depending on the scenario setup. The parameters are the number of stacked Transformer encoders and the dimension of their positional feedforward network block. We theorize that the depth and width of the communication channel create a trade-off between the ability to learn a complex joint policy (eg. kiting, hit-and-run), and the time it takes to learn an effective policy.
We perform a hyperparameter search for each scenario, for which detailed results are included in the Appendix, and use the values presented in Table 2.
Mixer compatibility
Both MACTAS and MAIC are agnostic to the value function decomposition algorithm, so we compare both algorithms in combination with commonly used mixers: VDN, QMIX, and QPLEX. Each combination named MACTAS+VDN, MAIC+QPLEX, etc., is tested across all maps and compared separately to provide a fair comparison and isolate the effect of the mixer algorithm.
4.3 Results
The basic results are presented in Fig. 3. We compare MACTAS with MAIC on 6 maps and 3 mixers, which makes 18 combinations. If we consider the area under the learning curve, MACTAS performs better than MAIC in 9 cases, MAIC performs better in 3 cases and there are 6 draws. If we consider the final performance, MACTAS yields better results 6 times, MAIC yields better results 3 times, and there are 9 draws.
On the 2c_vs_64zg map, all algorithms except MAIC+VDN achieve the same final performance. In terms of the area under the learning curve MACTAS performs similarly to MAIC for the QMIX and VDN mixers and better for the QPLEX mixer. MACTAS+QMIX and MAIC+QMIX give jointly the best results on this map. On the 5m_vs_6m map MACTAS+QPLEX achieves the same final result as MAIC+QPLEX but with a better area under the learning curve. MACTAS+QMIX gives the same final result as MAIC+QMIX, but has a worse area under the learning curve (MAIC+QMIX drops around two million steps). MACTAS+VDN achieves a better final result than MAIC+VDN with a similar area under the learning curve (MAIC+VDN starts to learn faster but is unable to go beyond $60\%$ of the mean of the won battles). On this map MACTAS+QPLEX gives the best result. We observe that on the 27m_vs_30m map, our algorithm achieves a better area under the learning curve for all three mixers, and for QPLEX, it also achieves a better final performance. For the other mixers, the final performance is similar. We hypothesize that the good performance of our algorithm is the case because the number of parameters of our algorithm is independent of the number of agents, and thus our method scales well in environments with a large number of agents. Overall, MACTAS+QMIX gives the best result on this map. On the MMM2 map, MAIC performs better for all three mixers in terms of the final result. However, the areas under the learning curves are similar and for MACTAS+QMIX and for MAIC+QMIX, as our algorithm starts to learn faster. MAIC+QPLEX performs the best on this map. This is the only map on which a version of our algorithm does not give the best result. On the 6h_vs_8z map MACTAS performs better than MAIC for the QMIX mixer in terms of the area under the learning curve and the final result. MACTAS+VDN gives a similar final result as MAIC+VDN, but with a better area under the learning curve. MACTAS+QPLEX achieves a similar final result as MAIC+QPLEX, but with a worse area under the learning curve. Overall, MACTAS+QMIX and MAIC+QPLEX give jointly the best result on this map. On the 3s5z_vs_3s6z map, MACTAS+QMIX and MACTAS+VDN give better results in terms of both the area under the learning curve and the final result than MAIC+QMIX and MAIC+VDN, respectively. MACTAS+QPLEX gives a similar final result to MAIC+QPLEX, but with a worse area under the learning curve. MACTAS+VDN performs the best on this map.
4.4 Ablations
There are two objectives of the ablation study. Firstly, we notice that the Transformer module has residual connections inside it. The question is whether the residual connection around it (the arrow between GRU and “++” in Fig. 2) is helpful indeed. Figure 5 shows the impact of the residual connection on MACTAS+QMIX. We observe that the residual connection is crucial for the 3s5z_vs_3s6z and 6h_vs_8z maps.
Secondly, we verify the role of our proposed exploration scheme. Figure 4 shows the impact of our exploration scheme for MACTAS+QMIX. We observe that our exploration method improves the results on the 3s5z_vs_3s6z map and is indifferent on the 27m_vs_30m and MMM2 maps. We used softmax temperature $0.33$ for the 3s5z_vs_3s6z map and $0.25$ for the 27m_vs_30m and MMM2 maps.
4.5 Discussion
[…]
The modular nature of MACTAS is its great advantage. There are widespread variants of the Transformer code optimized for different machines. This makes the implementation and deployment of MACTAS in a given multi-agent system straightforward.
Limitations
A moderate sensitivity of the performance of our proposed communication module to its shape requires searching for its best setting in additional runs. This slows the deployment of MACTAS.
5 Conclusions and Future Work
[…]
We observe that $\epsilon$-greedy exploration applied independently by agents in MARL leaves room for improvement since it condones the agents’ policies falling into suboptimal Nash equilibria. We addressed this issue by proposing a combination of $\epsilon$-greedy and Boltzmann exploration, which increased the performance of MACTAS+QMIX on the 3s5z_vs_3s6z SMAC map. However, we believe that this issue requires further research and intend to undertake it in the future.
Acknowledgements
We gratefully acknowledge the Polish high-performance computing infrastructure PLGrid (HPC Center: ACK Cyfronet AGH) for providing computer facilities and support within the computational grant no. PLG/2025/018560 and no. PLG/2025/017992. We thank the University of Warsaw for providing access to the computing infrastructure.
Appendix A StarCraft Multi-Agent Challenge (SMAC)
As a benchmark for MACTAS, we chose the StarCraft Multi-Agent Challenge (SMAC) (Samvelyan et al. 2019), which is a simplified version of the StarCraft II game. The official repository is available on https://github.com/oxwhirl/smac. SMAC uses PySc2 to communicate with the game engine https://github.com/google-deepmind/pysc2. SMAC is characterized by decentralized management of each unit, and each unit is an independent agent. This feature allows for the testing of decentralized algorithms, including those originating from MARL. Benchmark scenarios are limited to combat with a limited number of units. Those scenarios do not account for the economic part of the game, the so-called “macro” part of the game – collecting resources and building structures. Resources in the game (gas and minerals) are required to build units and structures. Structures allow for the recruitment of units and their upgrading. The considered scenarios do not involve “worker” units, which collect resources. Instead, the scenarios focus on ’micro,’ the term used in the StarCraft II community for unit control during combat. The challenge for a human player is multitasking between these two tasks – developing the game economy and controlling the fighting units. SMAC focuses only on combat, thereby ensuring this benchmark repeatability, as each iteration involves a fight between the same specification units, eliminating the uncertainty with unit upgrades and different strategies for the economic part of the game. In StarCraft II, there are three races, and our experiments covered controlling units of all of them. The way to run SMAC is described in our repository.
Appendix B Detailed Results
Figures 6, 7, and 8 contain more detailed results of our experiments described in Section 4.3. The curves on the plots represent the mean percentage test game win rate during the algorithm training for 6 different environments and various mixers. The shaded area around the curves is the standard deviation of the results among 5 different seeds (1 to 5) of the random number generator. For each seed, the mean percentage test win rate was calculated based on 32 independent test episodes. We observe robustness of MACTAS with the QMIX mixer – on the most challenging environments 6h_vs_8z and 3s5z_vs_3s6z, it outperforms MAIC+QMIX by a large margin; on 27m_vs_30m, it learns much faster; on 2c_vs_64zg and 3s5z_vs_3s6z, it performs similarly, and only slightly worse on the MMM2 map. For the QPLEX mixer, our method performs much better than MAIC on the 2c_vs_64zg and 27m_vs_30m maps, slightly better on the 5m_vs_6m, slightly worse on MMM2, learns slower but achieves similar final result on the 6h_vs_8z and 3s5z_vs_3s6z maps. Our method combined with the VDN mixer gives significantly better results than MAIC on the 3s5z_vs_3s6z map, slightly better on 6h_vs_8z, 27m_vs_30m, 5m_vs_6m, marginally better on the 2c_vs_64zg map and marginally worse on the MMM2 map.
Figure 6: Means and standard deviations of the percentage wins in the test games during training for QMIX with MAIC, MACTAS, and bare.
Figure 7: Means and standard deviations of the percentage wins in the test games during training for QPLEX with MAIC, MACTAS, and bare.
Figure 8: Means and standard deviations of the percentage wins in the test games during training for VDN with MAIC, MACTAS, and bare.
Appendix C Tactics used by MACTAS
The SMAC challenge forces agents to learn various tactics, such as attacking from distance melee units (’kiting’) or using movement advantage (’hit and run’). In this section, we provide an overview of the tactics used by our model based on our understanding of StarCraft II. We observed differences in the behavior of our model depending on the random seed. The description of tactics is based on replays of our models for seed = 1.
C.1 2c_vs_64zg
Figure 9: 2c_vs_64zg replay.
In this scenario, agents control 2 colossi units. Colossi can move through hills between the high and the low ground. Against them, 64 enemy zerglings are deployed, which are fast ground units. Zerglings are most effective when they surround an opposite unit. In the scenario 2c_vs_64zg, the colossi start on the hill, while the zerglings are split into two groups of 32 on the left and the right, both of them on the low ground. Figure 9 presents the following tactics: colossi attack the right group of zerglings and wait for the left group to climb the hill. Then, colossi use the ability to traverse the hill and attack the remaining zerglings, which are forced to run around the hill. In this way, the colossi avoid being surrounded and thereby defeated.
C.2 27m_vs_30m and 5m_vs_6m
Figure 10: 27m_vs_30m replay.
Figure 11: 5m_vs_6m replay.
The 27m_vs_30m and the 5m_vs_6m scenarios consist of marine units only. They are similar in the way that the opposing force numerically outnumbers the agents under control. In Figure 10, the agent’s units place themselves in a semicircle formation, also known as flanking, which maximizes the number of shooting units, while the enemies block each other. Units controlled by MACTAS focus fire on enemy units with lower health, in order to quickly reduce the number of opponents. Furthermore, units with low health retreat and continue fighting, while avoiding being shot. This behavior is presented in Figure 11: Agents with low health retreat when others focus fire on one enemy unit. The winning strategy is similar for both 5m_vs_6m and 27m_vs_30m. 5m_vs_6m is labeled as a hard benchmark, while 27m_vs_30m is classified as super-hard. The percentage imbalance between units is higher in 5m_vs_6m. However, 27m_vs_30m has a larger number of units that need to cooperate.
C.3 MMM2
Figure 12: MMM2 replay.
The MMM2 is a scenario with a diverse set of units containing 1 medivac (medic), 2 marauders, and 7 marines fighting a team with 1 more marauder and 1 more marine. Medivac cannot attack, but can heal units as long as it has energy. MACTAS ground units focus fire on enemy units and eliminate them one by one, and fall back when their health is low. The enemy medivac does not change healed unit, while the MACTAS’ medivac changes healed unit to the one with a critically low health or to the currently attacked unit. Units trained with MACTAS move medivac to the front line and focus enemy fire on it, and when its health level is low, retreat it. MACTAS’ medivac taking enemy fire secures outnumbered marines and marauders. Typically, human players try to save their medivacs because they are used for healing units between fights during a regular game. Only under certain conditions do they sacrifice their medivacs; for example, when they are breaking through an opponent’s line of defense to focus enemy fire on the medivacs and away from damage-dealing units. In this scenario, we can observe similar behavior in the MACTAS medivac control, which results in winning combat against an opponent who has a numerical advantage.
C.4 6h_vs_8z
Figure 13: 6h_vs_8z replay.
Agents control 6 range shooting units, hydralisks, while the enemy controls 8 melee zealots. In Figure 13, we can see a tactic called “kiting” – hydralisks split into two groups and are attacking zealots from distances. Enemy zealots attack the nearest unit – our agents learned to move enough to change the focus of approaching units and then continue striking. Notably, the zealot units within this scenario are not upgraded with the “charge” ability, which would otherwise allow them to have a significant increase in movement speed and easily reach hydralisks.
C.5 3s5z_vs_3s6z
Figure 14: 3s5z_vs_3s6z replay.
MACTAS controls 3 stalkers and 5 zealots, while the enemy has one more zealot. The stalkers are fast, range-attacking units, but with weak defense. On the other hand, zealots have good defense, but need to get close to the opponent unit. The tactic in this scenario is interesting. Using one stalker as a decoy, MACTAS first defeats the enemy stalkers and then proceeds to eliminate the rest of the enemy force. One agent, a stalker, sacrifices himself and focuses a few enemy zealots on himself. In Figure 14, we can see 5 enemy zealots following the stalker marked as “Local Player”. The remaining MACTAS’ 2 stalkers and 5 zealots attack 3 enemy stalkers and 1 zealot with an advantage. This move results in the elimination of enemy stalkers and the zealot. In the end, MACTAS’ units eliminate zealots lured by the sacrificed stalker.
Appendix D Sensitivity analysis
Figure 15: MACTAS; sensitivity of the results to architecture size with the QMIX mixer.
Figure 16: MACTAS; sensitivity of the results to architecture size with the QPLEX mixer.
Figure 17: MACTAS; sensitivity of the results to architecture size with the VDN mixer. Due to the technical issues, the version with 3 Transformer layers and a width of the Transformer feedforward layer equal to 512 was evaluated on 4 seeds of the random generator (1 to 4).
Figure 18: MACTAS; sensitivity of the results to the dropout parameter in the Transformer block. The default value of this parameter is $10\%$.
Figure 19: MACTAS; sensitivity of the top 2 exploration method to the temperature coefficient.
We evaluated the choice of the number of Transformer encoder layers and the width of the Transformer feedforward layer for MACTAS with QMIX, QPLEX, and VDN as a mixer in five environments: 2c_vs_64zg, 5m_vs_6m, 27m_vs_30m, MMM2 and 6h_vs_8z. We experimented with the dropout of the Transformer encoder on all six maps. We also assessed the sensitivity of our new top $k$ exploration technique to the softmax temperature in the case of MACTAS+QMIX and $k=2$ on three maps: 27m_vs_30m, MMM2 and 3s5z_vs_3s6z.
[…]
Figure 18 shows how the dropout parameter affects the performance of MACTAS. We observe that the absence of dropout harms the performance; however, the choice of the dropout fraction is not crucial when set in a reasonable interval; relevant differences can be observed only on the 3s5z_vs_3s6z map.
Figure 19 shows the impact of the softmax temperature parameter on the proposed exploration scheme. We observe that this method is sensitive to this parameter. For the 27m_vs_30m and MMM2 maps, the low temperatures work better, with $0.25$ achieving similar results to the $\epsilon$-greedy scheme, which is equivalent to choosing 0 as the temperature. In these cases, the new exploration scheme does not outperform the $\epsilon$-greedy exploration. We can observe an interesting situation on the 3s5z_vs_3s6z map. The top 2 exploration scheme outperforms the $\epsilon$-greedy exploration for the $0.33$ temperature. However, a temperature equal to $4$ gives a similar result to the $\epsilon$-greedy exploration. The in-between temperatures perform worse than both of these values, showing that there is no monotonic relationship between the temperature and the performance.
Appendix E Software
Table 3 contains a list of key programming libraries used during our experiments. The complete list of the programming libraries is available in the MACTAS’ official code repository in file requirements.txt.
Library name Version
Python 3.9.21 or 3.9.22
numpy 1.23.1
nvidia-cublas-cu12 12.4.5.8
nvidia-cuda-cupti-cu12 12.4.127
nvidia-cuda-nvrtc-cu12 12.4.127
nvidia-cuda-runtime-cu12 12.4.127
nvidia-cudnn-cu12 9.1.0.70
nvidia-cufft-cu12 11.2.1.3
nvidia-curand-cu12 10.3.5.147
nvidia-cusolver-cu12 11.6.1.9
nvidia-cusparse-cu12 12.3.1.170
nvidia-nccl-cu12 2.21.5
nvidia-nvjitlink-cu12 12.4.127
nvidia-nvtx-cu12 12.4.127
scipy 1.13.1
torch 2.5.1
Table 3: Essential libraries in programming environment.
Appendix F Hardware
Cluster name GPU device GPU driver version CPU device Operating system Total RAM
Cluster 1 NVIDIA A100 SXM4 80GB 535.154.05 AMD EPYC 7742 64-Core Processor Linux-5.15.0-1059-nvidia-x86_64-with-glibc2.35 1024 GB
Cluster 2 NVIDIA A100 SXM4 40GB 570.133.20 AMD EPYC 7742 64-Core Processor Linux-5.14.0-503.26.1.el9_5.x86_64-x86_64-with-glibc2.34 1024 GB
Cluster 3 NVIDIA A100 PCIe 80GB 565.57.01 AMD EPYC 7713 64-Core Processor Linux-6.8.0-64-generic-x86_64-with-glibc2.39 2048 GB
Table 4: Specification of the computational clusters used in the experiments.
Here we present the hardware used in our experiments. We used NVIDIA Ampere GPU architecture on three computational clusters. The specific setup is presented in Table 4.

### v2

@@
 MACTAS: Self-Attention-Based Module for Inter-Agent Communication in Multi-Agent Reinforcement Learning
-Abstract
+Abstract.
 Communication is essential for the collective execution of complex tasks by human agents, motivating interest in communication mechanisms for multi-agent reinforcement learning (MARL).
@@
 Notably, it includes a fixed number of trainable parameters, independent of the number of agents.
-Experimental results on the SMAC benchmark demonstrate the effectiveness of our approach, which achieves state-of-the-art performance on several maps.
-1 Introduction
-Multi-agent reinforcement learning (MARL) (Ning and Xie 2024; Du and Ding 2021; Nguyen, Nguyen, and Nahavandi 2020; Oroojlooy and Hajinezhad 2023; Wong et al.
-2023) promises efficient control in large-scale systems through the collective learning of many agents in dynamic environments.
+Experimental results on the SMAC and SMACv2 benchmarks demonstrate the effectiveness of our approach, which achieves state-of-the-art performance on a number of maps.
+Key words and phrases:
+Multi-Agent Reinforcement Learning, Cooperation in Multi-Agent Reinforcement Learning, Communication in Multi-Agent Reinforcement Learning
+1. Introduction
+Multi-agent reinforcement learning (MARL) Ning and Xie (2024); Du and Ding (2021); Nguyen et al.
+(2020); Oroojlooy and Hajinezhad (2023); Wong et al.
+(2023) promises efficient control in large-scale systems through the collective learning of many agents in dynamic environments.
 Multiple agents interact within a shared environment to learn optimal policies.
@@
 The challenges and complexities in using MARL for controlling multi-agent systems include coordination among agents, non-stationary environments due to evolving agent policies, and scalability issues.
-Advanced MARL approaches often involve sophisticated techniques such as decentralized training, communication protocols among agents, and hierarchical methods to manage multi-agent interactions as effectively as possible.
-Theoretical advancements and practical applications of MARL demonstrate its potential in diverse fields ranging from robotics (Orr and Dutta 2023) and autonomous driving (Yadav, Mishra, and Kim 2023; Schmidt et al.
-2022; Althamary, Huang, and Lin 2019) to telecommunications (Li et al.
-2022b; Feriani and Hossain 2021).
-Figure 1: 3s5z_vs_3s6z scenario in the StarCraft Multi-Agent Challenge (SMAC).
-The most prominent research topics in MARL are the action-value function decompositions (Wang et al.
-2021; Liu, Zhu, and Chen 2023), where agents share a joint reward according to their contribution, and inter-agent communication (Liu et al.
-2023; Guo, Shi, and Fan 2023), where agents exchange information in order to coordinate their joint policy.
-The most challenging multi-agent environments are partially observable, usually via limiting agents’ information to their close proximity, and the successful achievement of their common goals is predicated upon coordination, which needs to involve some form of communication.
+Advanced MARL approaches often involve sophisticated techniques, including decentralized training, communication protocols among agents, and hierarchical methods to manage multi-agent interactions as effectively as possible.
+Theoretical advancements and practical applications of MARL demonstrate its potential in diverse fields ranging from robotics Orr and Dutta (2023) and autonomous driving Yadav et al.
+(2023); Schmidt et al.
+(2022); Althamary et al.
+(2019) to telecommunications Li et al.
+(2022b); Feriani and Hossain (2021).
+Figure 1. 3s5z_vs_3s6z scenario in the StarCraft Multi-Agent Challenge (SMAC).
+The most prominent research topics in MARL are the action-value function decompositions Wang et al.
+(2021); Liu et al.
+(2023b), where agents share a joint reward according to their contributions, and inter-agent communication Liu et al.
+(2023a); Guo et al.
+(2023), where agents exchange information in order to coordinate their joint policy.
+The most challenging multi-agent environments are partially observable, typically by limiting agents’ information to their close proximity, and the achievement of their common goals is predicated upon coordination, which needs to involve some form of communication.
 In most research, the action-value function and inter-agent communication are issues addressed separately.
@@
 Moreover, in most challenging MARL tasks, some form of inter-agent communication is a necessity, which can be provided relatively easily (e.g., through WiFi).
-Therefore, we introduce a communication protocol based on collecting information from each agent in a central entity and broadcasting the combined information to all agents.
-Because information aggregation is based on attention, our proposed architecture requires a constant number of weights plus an embedding for each agent.
+Therefore, we introduce a communication protocol that collects information from each agent in a central entity and broadcasts the combined information to all agents.
+Because information aggregation is based on attention, our proposed architecture requires a constant number of weights, plus an embedding, for each agent.
 It is easy to apply using readily available and efficient implementations of the Transformer encoder layer.
@@
   • We introduce Multi-Agent Communication via Transformer on Agents’ States (MACTAS) module for inter-agent communication in MARL. This module allows the agent to share their state of consciousness.
-  • We verify MACTAS experimentally on several challenging maps from SMAC, giving excellent results and establishing state-of-the-art in some cases.
-2 Related Work
-Single-Agent Reinforcement Learning (SARL) (Sutton and Barto 2018) addresses trial-and-error learning of sequential decision making under uncertainty.
-In Multi-Agent Reinforcement Learning (MARL) (Ning and Xie 2024), a group of agents learns to act in the same dynamic environment, collectively modifying its state.
+  • We verify MACTAS experimentally on several challenging maps from SMACSamvelyan et al.
+(2019) and SMACv2 Ellis et al.
+(2023), giving excellent results and establishing state-of-the-art in some cases.
+2. Related Work
+Single-Agent Reinforcement Learning (SARL) Sutton and Barto (2018) addresses trial-and-error learning in sequential decision-making under uncertainty.
+In Multi-Agent Reinforcement Learning (MARL) Ning and Xie (2024), a group of agents learns to act in the same dynamic environment, collectively modifying its state.
 There are various settings for MARL, depending on whether the agents operate synchronously or asynchronously, whether they have common or adverse goals, and other circumstances.
 In this paper, we focus on the synchronous and cooperative MARL, with a single reward value for all agents at every time instance.
-MARL has been addressed by extending well-established SARL algorithms: Multi-Agent SAC (Wei et al. 2018), Multi-Agent PPO (Yu et al. 2022), Multi-Agent DDPG (Lowe et al. 2017).
-Generally, cooperative MARL adopts a centralized training with a decentralized execution (CTDE) paradigm, which suffers from the global action-value function, whose complexity grows exponentially with the number of agents (Kraemer and Banerjee 2016).
+MARL has been addressed by extending well-established SARL algorithms, including Multi-Agent SAC Wei et al. (2018), Multi-Agent PPO Yu et al. (2022), and Multi-Agent DDPG Lowe et al. (2017).
+Methods of cooperative MARL often adopt a centralized training with decentralized execution (CTDE) paradigm, in which agents share information during training but act based solely on their own observations during execution.
+However, this paradigm suffers from the global action-value function, whose complexity grows exponentially with the number of agents Kraemer and Banerjee (2016).
 Therefore, action-value function decomposition is a fundamental problem in MARL.
@@
 These modules are fed with agents’ observations and their previous actions and produce agents’ contributions to the architecture’s output.
-The best action for the agent is the one that produces the highest contribution.
-In VDN (Sunehag et al.
-2018), these contributions are just added.
-In QMIX (Rashid et al.
-2018), a neural module, the mixing network, produces its weighted sum with positive weights.
-In Weighted QMIX (Rashid et al.
-2020), the architecture is trained with a special weight assigned to the collective actions deemed the best.
-In QTRAN (Son et al.
-2019), an additional module fed with all the actions of all agents corrects the weighted sum of the agents’ contributions.
-In Qatten (Yang et al.
-2020), the mixing network is based on attention.
-In CDS (Li et al.
-2021), a component of the mixing network optimizes the amount of exploration undertaken by the agents.
-In DVD (Li et al.
-2022a), the internal state of the modules assigned to agents is also passed to the mixing networks.
-In SHAQ (Wang et al.
-2022), the mixing network is enriched by several components based on Shapley value theory.
-In QPLEX (Wang et al.
-2021), a carefully engineered duplex dueling of pairs of contributions of agents is introduced into the mixing network.
-In NA2Q (Liu, Zhu, and Chen 2023), a similar duplex dueling is performed with the use of attention.
-Another fundamental problem in the cooperative MARL is the coordination of agents’ activity with a certain communication protocol.
-A number of solutions have been proposed for that purpose, including CommNet (Sukhbaatar, Szlam, and Fergus 2016) which introduces differentiable communication by aggregating a learned broadcast vector from each agent, TarMAC (Das et al.
-2019) uses signature-based soft attention to direct agent communication towards recipients, NDQ (Wang et al.
-2019) applies two forms of regularization – mutual information maximization and message entropy minimization to optimize message content, GA-Comm (Liu et al.
-2020) uses two-stage attention mechanism to limit number of messages exchanged between agents, IS (Kim, Park, and Sung 2021) encodes trajectory of future actions for each agent and shares them with others, in MAIC (Yuan et al.
-2022) agents explicitly model their teammates and use it to bias their value function, DHCG (Liu et al.
-2023) learns a dependency graph between agents and uses it to limit the number of communication edges, TEM (Guo, Shi, and Fan 2023) which uses a Transformer to simulate communication resembling email exchange, where agents only directly communicate with teammates nearby, but can forward messages from one to another, and ACE (Li et al.
-2023) which proposes bidirectional action-dependent Q-learning.
-In addition, there are solutions in which the agents communicate in natural language, including: TWOSOME (Tan et al.
-2024) which uses an LLM to evaluate the probabilities of joint action and applies LoRa for efficient PPO training, and Verco (Li et al.
-2024) which separates the output of the LLM into communication and actions, allowing LLM agents to exchange interpretable verbal messages.
-In our setting, we consider communication methods that can be connected with an arbitrary action-value decomposition method.
+The best action for the agent is the one that makes the highest contribution.
+In VDN Sunehag et al.
+(2018), these contributions are just added.
+In QMIX Rashid et al.
+(2018), a neural module, the mixing network, produces its weighted sum with positive weights.
+In Weighted QMIX Rashid et al.
+(2020), the architecture is trained with a special weight assigned to the collective actions deemed the best.
+In QTRAN Son et al.
+(2019), an additional module fed with all the actions of all agents corrects the weighted sum of the agents’ contributions.
+In Qatten Yang et al.
+(2020), the mixing network is based on attention.
+In CDS Li et al.
+(2021), a component of the mixing network optimizes the amount of exploration undertaken by the agents.
+In DVD Li et al.
+(2022a), the internal state of the modules assigned to agents is also passed to the mixing networks.
+In SHAQ Wang et al.
+(2022), the mixing network is enriched by several components based on Shapley value theory.
+In QPLEX Wang et al.
+(2021), a carefully engineered duplex dueling of pairs of contributions of agents is introduced into the mixing network.
+In NA2Q Liu et al.
+(2023b), a similar duplex dueling is performed with the use of attention.
+Another fundamental problem in cooperative MARL is the coordination of agents’ activity for which a certain communication protocol is required.
+A number of solutions have been proposed for this purpose: CommNet Sukhbaatar et al.
+(2016) introduces differentiable communication by aggregating a learned broadcast vector from each agent, TarMAC Das et al.
+(2019) uses signature-based soft attention to direct agent communication towards recipients, NDQ Wang et al.
+(2019) applies two forms of regularization – mutual information maximization and message entropy minimization to optimize message content, GA-Comm Liu et al.
+(2020) uses two-stage attention mechanism to limit number of messages exchanged between agents, IS Kim et al.
+(2021) encodes trajectory of future actions for each agent and shares them with others, in MAIC Yuan et al.
+(2022) agents explicitly model their teammates and use it to bias their value function, DHCG Liu et al.
+(2023a) learns a dependency graph between agents and uses it to limit the number of communication edges, TEM Guo et al.
+(2023) uses a Transformer to simulate communication resembling email exchange, where agents only directly communicate with teammates nearby, but can forward messages from one to another, ACE Li et al.
+(2023) proposes bidirectional action-dependent Q-learning.
+In addition, there are solutions in which the agents communicate in natural language, including: TWOSOME Tan et al.
+(2024) which uses a large language model (LLM) to evaluate the probabilities of joint action and applies LoRa for efficient PPO training, and Verco Li et al.
+(2024) which separates the output of the LLM into communication and actions, allowing LLM agents to exchange interpretable verbal messages.
+We acknowledge various applications of attention or Transformer architectures in MARL algorithms, such as in Qatten mixing network, or in communication algorithms like NA2Q, GA-Comm, and TEM.
+None of the previous works has applied the Transformer architecture directly as a communication protocol.
+Our work fills this gap.
+We consider communication methods that can be coupled with an arbitrary action-value decomposition method.
 We also allow agents to send messages based only on their local information (current observation, historical observations, and historical actions).
-In this framework, the best results so far have been achieved by MAIC (Yuan et al.
-2022).
-3 Method
-Figure 2: The proposed MACTAS architecture.
-3.1 Problem definition
-In this paper, we discuss the problem of MARL using the formalism of a cooperative Decentralized Partially Observable Markov Decision Process (DecPOMDP) with communication.
-A DecPOMDP is defined by a tuple, $\langle\mathbb{N},\mathbb{S},\mathbb{A},\mathcal{P},\mathcal{O},\mathcal{R}\rangle$, where $\mathbb{N}=\{1,\dots,n\}$ is a team of agents, $\mathbb{S}$ is a space of environmental states, $\mathbb{A}=\mathbb{A}^{1}\times\dots\times\mathbb{A}^{n}$ is a space of team actions of the agents, $\mathcal{P}$ is the state transition probability, $\mathcal{O}$ is the observation function, and $\mathcal{R}$ is the reward function.
-At time $t=1,2,\dots$ each $i$-th agent makes an observation, $o^{i}_{t}=\mathcal{O}^{i}(s_{t})$, of the environment state $s_{t}$.
+In this framework, the best results so far have been achieved by MAIC Yuan et al.
+(2022).
+3. Method
+Figure 2. The proposed MACTAS architecture.
+3.1. Problem definition
+In this paper, we discuss the problem of MARL using the formalism of a cooperative Decentralized Partially Observable Markov Decision Process (Dec-POMDP) with communication.
+Dec-POMDP is defined by a tuple $\langle\mathbb{N},\mathbb{S},\mathbb{A},\mathcal{R},\mathcal{P},\Omega,\mathcal{O},\gamma\rangle$, where $\mathbb{N}=\{1,\dots,n\}$ represents the team of agents, $\mathbb{S}$ is the global environmental state space, $\mathbb{A}=\mathbb{A}^{1}\times\dots\times\mathbb{A}^{n}$ is the joint action space, $\mathcal{R}$ is the reward function, $\mathcal{P}$ is the state transition probability function, $\Omega=\Omega^{1}\times\dots\times\Omega^{n}$ is the joint observation space, $\mathcal{O}$ is the observation function and $\gamma$ is the discount factor.
+At time $t=1,2,\dots$ each $i$-th agent makes an observation, $o^{i}_{t}=\mathcal{O}^{i}(s_{t}),o^{i}_{t}\in\Omega^{i}$, of the environment state $s_{t}\in\mathbb{S}$.
 Agents may exchange messages with others through a certain channel that may generally refine these messages with some processing.
@@
 Based on thereof and a policy, $\pi^{i}$, the agent performs an action, $a^{i}_{t}\in\mathbb{A}^{i}$, $a^{i}_{t}\sim\pi^{i}(\cdot|\mathcal{I}^{i}_{t})$.
-The team action $a_{t}=[(a^{1}_{t}),\dots,(a^{n}_{t})]$ impacts the next environment state $s_{t+1}\sim\mathcal{P}(\cdot|s_{t},a_{t})$ and the reward $r_{t}=\mathcal{R}(s_{t+1})$ is given collectively to all agents.
-The goal of the agents’ learning is to optimize their team policy $\pi=[\pi^{1},\dots,\pi^{n}]$ and the message exchange protocol for the team to expect in each environment state $s$ the highest sum of discounted rewards
-$\mathcal{V}^{\pi}(s)=E\left(\sum_{i\geq 0}\gamma^{i}r_{t+i}\Big{|}s_{t}=s,\text{policy in use}=\pi\right),$
-where $\gamma\in[0,1]$ is the discount factor.
-The goal of this paper is to establish a message exchange protocol that supports the above goal of learning.
-3.2 MACTAS
+The team action $a_{t}=[(a^{1}_{t}),\dots,(a^{n}_{t})]$ impacts the next environment state $s_{t+1}\sim\mathcal{P}(\cdot|s_{t},a_{t})$ and the reward $r_{t}=\mathcal{R}(s_{t},a_{t},s_{t+1})$ is given collectively to all agents.
+The goal of the agents’ learning is to optimize their team policy $\pi=[\pi^{1},\dots,\pi^{n}]$ to expect in each environment state $s$ the highest sum of discounted rewards
+$\mathcal{V}^{\pi}(s)=\mathbb{E}\left(\sum_{i\geq 0}\gamma^{i}r_{t+i}\Big|s_{t}=s,\text{policy in use}=\pi\right),$
+where $\gamma\in[0,1]$.
+The goal of this paper is to establish the message exchange protocol that supports the above goal of learning.
+3.2. MACTAS
 Our proposed Multi-Agent Communication via Transformer on Agents States (MACTAS) architecture is presented in Fig.
@@
 These modules are fed with previous actions of these agents, $a^{i}_{t-1}$, and their current observations, $o^{i}_{t}$.
-They produce agents’ contributions, $\widehat{Q}^{i}_{t}$, to future discounted rewards and perhaps some other vectors (Li et al.
-2022a).
+They produce agents’ contributions, $\widehat{Q}^{i}_{t}$, to future discounted rewards and perhaps some other vectors Li et al.
+(2022a).
 The outputs of the $f^{i}$ modules are fed to the mixing network that produces the value of the action-value function, $\widehat{Q}_{t}$.
@@
 In order to give a global view on the current state to the agent’s consciousness, we introduce a communication mechanism.
-We divide the $f^{i}$ module after its recurrent part and feed its state $h^{i}_{t}\in\mathbb{R}^{n_{h}}$ to the module $g$ where the states of consciousness of the agents are confronted.
-As a result, the $g$ module produces increments, $z^{i}_{t}\in\mathbb{R}^{n_{h}}$.
+We divide the $f^{i}$ module after its recurrent part and feed its state $h^{i}_{t}\in\real^{n_{h}}$ to the module $g$, where the states of consciousness of the agents are confronted.
+As a result, the $g$ module produces increments, $z^{i}_{t}\in\real^{n_{h}}$.
 The sum $z^{i}_{t}+h^{i}_{t}$ is then passed to the second MLP part of the $f^{i}$ module.
@@
 The minimal requirement for the $g$ module is as follows:
-  1. 1.
+  1. (1)
     It is fed with $n$ vectors and produces $n$ vectors, both of size $\dim(h^{i}_{t})$. $i$-th input vector corresponds to the $i$-th output vector.
-  2. 2.
-    It is a layered structure and the layers include self-attention (Vaswani et al. 2017).
-Introducing $g$ satisfying the above requirements in their basic form is enough to outperform MARL without communication by a large margin.
-However, what follows from our experiments is that usually the best performing form of the $g$ module is a stack of Transformer (encoder) modules (Vaswani et al.
-2017), with the following settings:
-  1. 1.
+  2. (2)
+    It is a layered structure and the layers include self-attention Vaswani et al. (2017).
+Introducing $g$ that satisfies the above requirements in their basic form is enough to outperform MARL without communication by a large margin.
+However, what follows from our experiments is that usually the best performing form of the $g$ module is a stack of Transformer (encoder) modules Vaswani et al.
+(2017), with the following settings:
+  1. (1)
     There are no positional embeddings. The agents’ embeddings are fed to $f^{i}$ instead.
-  2. 2.
+  2. (2)
     There is no masking.
-  3. 3.
-    We use a separate Adam optimizer (Kingma and Ba 2017) for the Transformer parameters; the rest of the architecture is trained with its originally proposed optimizer (typically it is RMSProp: Tieleman and Hinton 2012).
-  4. 4.
-    The weights of the output layer of the top Transformer module are initialized with zeros. Therefore, initially, this module simply passes the input forward.
-Good performance of the Transformer in this role is an expected result.
+  3. (3)
+    We use a separate Adam optimizer Kingma and Ba (2017) for the Transformer parameters; the rest of the architecture is trained with its originally proposed optimizer (typically it is RMSProp Tieleman and Hinton (2012)).
+  4. (4)
+    The weights of the output layer of the top Transformer module are initialized with zeros. Therefore, initially, this module does not impact the state of consciousness of the agents.
+A good performance of the Transformer in this role is an expected result.
 This architecture has been specifically optimized to transform a set of embeddings into another set of the same size, which is exactly what we require in the $g$ module.
-3.3 Implementation considerations
+3.3. Implementation considerations
 The basic approach to implement a multi-agent system with our proposed MACTAS module is to set up the Intermediate Processing Unit (IPU) on which the module is run.
@@
 Usually in multi-agent systems, reducing the communication capacity to $O(n)$ and offloading the communication processing burden to the IPU is very convenient.
-In these systems, the agents are naturally connected through a network and their computational capacity is limited (e.g., they are mobile and battery powered), contrary to that of the network infrastructure.
+In these systems, the agents are naturally connected through a network and their computational capacity is limited (e.g., they are mobile and battery-powered), contrary to that of the network infrastructure.
 Still, these are the agents that need to take on the processing of their observations and control.
-In most inter-agent communication protocols, including MAIC (Yuan et al.
-2022), all pairs of agents need to be able to communicate, which translates into $O(n^{2})$ messages.
+In most inter-agent communication protocols, including MAIC Yuan et al.
+(2022), all pairs of agents need to be able to communicate, which translates into $O(n^{2})$ messages.
 Technically, the parts of the agents that communicate with others could be moved to an IPU.
-That would require $O(n)$ messages to and from the IPU.
+Communication through the IPU would require $O(n)$ messages from the agents to the IPU and back.
 Still, natural parallelization of MACTAS is its noticeable advantage.
@@
 In training, the agents need to aggregate increments in their shared weights.
-In the distributed implementation, it may be the case that some of the agents are out of reach of others and do not exchange information with them.
+In the distributed implementation, it may happen that some of the agents are out of reach of others and do not exchange information with them.
 That only means that self-attention in MACTAS is forced to assign zero weight to unavailable transmissions from those unheard agents.
-3.4 Exploration
+3.4. Exploration
 Learning in a multi-agent system with decentralized execution is hampered by agents’ policies falling into suboptimal Nash equilibria.
@@
 On average, only $n\epsilon$ agents explore.
-In order to learn that a certain combination of agents’ actions is better than the current Nash equilibrium, at least two agents need to explore and try a better combination.
-However, this is unlikely.
-What is more likely is that only one agent explores, which leads to a combination worse than greedy.
+To learn that a certain combination of agents’ actions is better than the current Nash equilibrium, at least two agents need to explore and try a better combination.
+However, this happens rarely.
+What happens more often is that only one agent explores, which leads to a combination worse than greedy.
 Note that this effect may be even stronger for a better approximation of the $Q$ function, as it leaves less room for accidental switches of the greedy action combinations.
@@
 Using $k=1$ is equivalent to the pure $\epsilon$-greedy policy.
-Selecting a small $k$ leads to exploring the region that is quite good according to the agent.
+Selecting a small $k$ leads to exploring a region that is quite good according to the agent.
 We observe that $k=2$ works well in practice.
 The temperature of the Boltzmann exploration controls the amount of extra exploration that we want to add.
-We plan to extend this idea and other exploration strategies for MARL in future work.
 Hyperparameter Value
@@
 Learning rate 0.0005
-Table 1: Basic experimental settings and hyperparameters. 1 for 2.5 mln. step maps. 2 for 7.5 mln. step maps.
+Table 1. Basic experimental settings and hyperparameters. 1 for 2.5 mln. step maps. 2 for 7.5 mln. step maps.
 Scenario Stacked encoders FFN dimension
@@
 3s5z_vs_3s6z 3 512
-Table 2: MACTAS shape for each scenario.
-Figure 3: Percentage of wins in test games in training time for the connection protocols MACTAS and MAIC, and the $Q$ architectures QMIX, QPLEX, and VDN.
-Figure 4: The impact of our exploration scheme on MACTAS+QMIX.
-Figure 5: The impact of our residual connection on MACTAS+QMIX.
-4 Experiments
-In this section, we present the results of our proposed MACTAS communication protocol in the StarCraft Multi-Agent Challenge (SMAC: Samvelyan et al.
-2019) environment.
-We compare it with the state-of-the-art MAIC (Yuan et al.
-2022), in combinations with 3 mixer algorithms: QMIX, QPLEX, VDN, and either MAIC or MACTAS communication protocol, on 6 maps.
+All SMACv2 scenarios 3 512
+Table 2. MACTAS shape for each scenario.
+Figure 3. Percentage of wins in test games in training time for the connection protocols MACTAS and MAIC, and the $Q$ architectures QMIX, QPLEX, and VDN.
+Figure 4. Percentage of wins and standard deviations in test games in training time for the connection protocols MACTAS, MAIC, and the bare mixer for the $Q$ architecture QMIX
+Figure 5. The impact of our exploration scheme on MACTAS+QMIX.
+Figure 6. The impact of our residual connection on MACTAS+QMIX.
+4. Experiments
+In this section, we present the results of our proposed MACTAS communication protocol in the StarCraft Multi-Agent Challenge environments (SMAC Samvelyan et al.
+(2019) and SMACv2 Ellis et al.
+(2023)).
+We compare it with the state-of-the-art MAIC Yuan et al.
+(2022), in combinations with 3 mixer algorithms: QMIX, QPLEX, and VDN, on 6 maps.
 We present results of two ablation studies regarding the design of neural network connectivity and the exploration form.
-Additionally, in the Appendix, we report the sensitivity study of MACTAS to hyperparameters, lower-level SMAC settings, and analyses of the most interesting strategies that have emerged in training.
-4.1 StarCraft Multi-Agent Challenge
-StarCraft Multi-Agent Challenge (SMAC) is a micromanagement benchmark where two opposite teams face each other in different scenarios (maps), one team being controlled by the RL agent and the other by a built-in algorithm.
+Additionally, in the Appendix, we report the sensitivity study of MACTAS to hyperparameters, lower-level SMAC and SMACv2 settings, and analyses of the most interesting strategies that have emerged in training.
+4.1. StarCraft Multi-Agent Challenge
+The StarCraft Multi-Agent Challenge (SMAC and SMACv2) is a micromanagement benchmark where two opposing teams face each other in different scenarios (maps), with one team controlled by the RL agent and the other by a built-in algorithm.
 Maps differ by difficulty, mainly due to an imbalance in the number of controlled units.
-For our experiments, we selected 2 hard and 4 super-hard scenarios, which are widely benchmarked in the MARL community:
+For our experiments, we selected 2 hard and 4 super-hard SMAC scenarios, which are widely benchmarked in the MARL community:
   • 2c_vs_64zg with 2 colossus units leveraging terrain advantage against 64 small enemy zerglings. Level: hard.
@@
   • 27m_vs_30m big battle of 27 marines conquering 30 marines. Level: super-hard.
-  • MMM2 scenario with a diverse units containing 1 medivac (medic), 2 marauders, and 7 marines fighting team with 1 more marauder and 1 more marine. Level: super-hard.
+  • MMM2 scenario with diverse units containing 1 medivac (medic), 2 marauders, and 7 marines fighting team with 1 more marauder and 1 more marine. Level: super-hard.
   • 6h_vs_8z with 6 shooting hydralisks against 8 malee zealots. Level: super-hard.
   • 3s5z_vs_3s6z scenario with 3 stalkers and 5 zealots against 3 stalkers and 6 zealots. Level: super-hard.
-In order to verify versatility of MACTAS algorithm, we run experiments on a diverse set of hard and super-hard scenarios.
+To verify the versatility of the MACTAS algorithm, we run experiments on a diverse set of hard and super-hard scenarios.
 Every map provides the opportunity to find a unique winning strategy.
-The agent requires learning basic skills such as navigating the battlefield, managing health and shield, or focusing the fire.
-All considered scenarios are imbalanced in favor of the enemies; therefore, simple strategies are not sufficient to win the episode.
-Popular observed winning policies include crossing the obstacle on 2c_vs_64zg, flanking on 27m_vs_30m, ’kiting’ on 6h_vs_8z, ’hit and run’ on 3s5z_vs_3s6z (Samvelyan et al.
-2019).
-4.2 Experimental setting
-We borrow the basic settings from NA2Q (Liu, Zhu, and Chen 2023) and extend their code with the implementation of MACTAS and MAIC.
-The algorithms are trained using the Double Q-Learning (Hasselt, Guez, and Silver 2016) algorithm.
-Each experiment is run for a fixed number of training episode steps – 2.5 million for 2c_vs_64zg, 5m_vs_6m, MMM2, 27m_vs_30m environments and 7.5 million for 6h_vs_8z, 3s5z_vs_3s6z environments, interleaved with test episodes every 2000 steps.
+The agent requires learning basic skills, such as navigating the battlefield, managing health and shield, or focusing fire.
+All considered scenarios are imbalanced in favor of the enemies; therefore, simple strategies are insufficient to win the episode.
+Popular observed winning policies include crossing the obstacle on 2c_vs_64zg, flanking on 27m_vs_30m, ’kiting’ on 6h_vs_8z, ’hit and run’ on 3s5z_vs_3s6z Samvelyan et al.
+(2019).
+In addition to classical SMAC, we benchmark the algorithms in its newer iteration, SMACv2 Ellis et al.
+(2023), which introduces new maps focused on scenario randomization.
+Each of the new maps defines the number of units in each team and the race (the same for both teams), while the exact types and starting positions are randomized according to the configuration.
+We use the default configuration for each scenario recommended in the SMACv2 repository.
+Our experiments focus on more challenging configurations, where the player team is outnumbered by the enemy team controlled by in-game AI, with 6 tested scenarios: zerg_5_vs_5, zerg_10_vs_11, zerg_20_vs_23, protoss_5_vs_5, protoss_10_vs_11, and protoss_20_vs_23.
+4.2. Experimental setting
+We borrow the basic settings from NA2Q Liu et al.
+(2023b) and extend their code with the implementation of MACTAS and MAIC.
+The algorithms are trained using the Double Q-Learning Hasselt et al.
+(2016) algorithm.
+Each experiment is run for a fixed number of training episode steps – 2.5 million for 2c_vs_64zg, 5m_vs_6m, MMM2, 27m_vs_30m environments and 7.5 million for 6h_vs_8z, 3s5z_vs_3s6z and all SMACv2 environments, interleaved with test episodes every 2000 steps.
 Each experiment is repeated for seeds $1,\dots,5$.
-We report the average percentage of wins (and standard deviations in the Appendix) in test games.
+We report the average percentage of wins in test games, along with the standard deviations in the Appendix.
 All hyperparameters not strictly related to the communication algorithm are shared between maps, algorithms, and mixers and are presented in Table 1.
-For MAIC communication, we use the same hyperparameters as in the original paper.
+For MAIC communication, we employ the same hyperparameters as those used in the original paper.
 MACTAS hyperparameters
@@
 Each combination named MACTAS+VDN, MAIC+QPLEX, etc., is tested across all maps and compared separately to provide a fair comparison and isolate the effect of the mixer algorithm.
-4.3 Results
+4.3. Results
+SMAC
 The basic results are presented in Fig.
@@
 For the other mixers, the final performance is similar.
-We hypothesize that the good performance of our algorithm is the case because the number of parameters of our algorithm is independent of the number of agents, and thus our method scales well in environments with a large number of agents.
+We hypothesize that the good performance of our algorithm is because the number of parameters in our algorithm is independent of the number of agents.
+Thus, our method scales well in environments with a large number of agents.
 Overall, MACTAS+QMIX gives the best result on this map.
@@
 MACTAS+VDN performs the best on this map.
-4.4 Ablations
+SMACv2
+The basic results are presented in Fig.
+4.
+We compare MACTAS with MAIC and the bare mixer on 6 maps and the QMIX mixer.
+In all scenarios, our algorithm outperforms MAIC and the bare mixer.
+We observe that the differences between MACTAS+QMIX and MAIC+QMIX are similar to the differences between MAIC+QMIX and the bare mixer.
+4.4. Ablations
 There are two objectives of the ablation study.
@@
 Figure 5 shows the impact of our exploration scheme for MACTAS+QMIX.
-We observe that our exploration method improves the results on the 3s5z_vs_3s6z map and is indifferent on the 27m_vs_30m and MMM2 maps.
-We used softmax temperature $0.33$ for the 3s5z_vs_3s6z map and $0.25$ for the 27m_vs_30m and MMM2 maps.
-4.5 Discussion
+We observe that our exploration method improves the results on the most challenging SMAC maps, which require $7.5$ million training steps: 3s5z_vs_3s6z, and 6h_8z.
+It is also indifferent on the maps: 2c_vs_64zg, 27m_vs_30m, and MMM2.
+Our exploration scheme slows down the training for the 5m_vs_6m map.
+We used softmax temperature $0.33$ for the map 3s5z_vs_3s6z and $0.25$ for the maps: 2c_vs_64zg, 5m_vs_6m, 27m_vs_30m, MMM2, and 6h_8z.
+4.5. Discussion
 […]
@@
 Limitations
-A moderate sensitivity of the performance of our proposed communication module to its shape requires searching for its best setting in additional runs. This slows the deployment of MACTAS.
-5 Conclusions and Future Work
+A moderate sensitivity of the performance of our proposed communication module to its shape requires searching for its best setting in additional runs.
+This slightly slows down the deployment of MACTAS.
+5. Conclusions and Future Work
 […]
 We observe that $\epsilon$-greedy exploration applied independently by agents in MARL leaves room for improvement since it condones the agents’ policies falling into suboptimal Nash equilibria.
-We addressed this issue by proposing a combination of $\epsilon$-greedy and Boltzmann exploration, which increased the performance of MACTAS+QMIX on the 3s5z_vs_3s6z SMAC map.
+We addressed this issue by proposing a combination of $\epsilon$-greedy and Boltzmann exploration, which increased the performance of MACTAS+QMIX on the 6h_vs_8z, and the 3s5z_vs_3s6z SMAC maps.
 However, we believe that this issue requires further research and intend to undertake it in the future.
-Acknowledgements
+{acks}
 We gratefully acknowledge the Polish high-performance computing infrastructure PLGrid (HPC Center: ACK Cyfronet AGH) for providing computer facilities and support within the computational grant no.
@@
 We thank the University of Warsaw for providing access to the computing infrastructure.
-Appendix A StarCraft Multi-Agent Challenge (SMAC)
-As a benchmark for MACTAS, we chose the StarCraft Multi-Agent Challenge (SMAC) (Samvelyan et al.
-2019), which is a simplified version of the StarCraft II game.
-The official repository is available on https://github.com/oxwhirl/smac.
+Appendix A StarCraft Multi-Agent Challenge (SMAC & SMACv2)
+As a benchmark for MACTAS, we chose the StarCraft Multi-Agent Challenge (SMAC) Samvelyan et al.
+(2019) and its newer version SMACv2 Ellis et al.
+(2023), which are simplified versions of the StarCraft II game.
+The framework exists in two versions – SMAC (code available at https://github.com/oxwhirl/smac), which introduced predefined deterministic scenarios that often include purposeful selection of units or obstacles placement, and SMACv2 (code available at https://github.com/oxwhirl/smacv2), which added new randomized scenarios based on plain maps and randomized units generation.
+From here onward, we will use ”SMACv1” to refer to the first version of the benchmark and ”SMAC” to refer to both SMACv1 and SMACv2, making an explicit distinction where necessary.
 SMAC uses PySc2 to communicate with the game engine https://github.com/google-deepmind/pysc2.
-SMAC is characterized by decentralized management of each unit, and each unit is an independent agent.
+SMAC is characterized by the decentralized management of each unit, where each unit acts as an independent agent.
 This feature allows for the testing of decentralized algorithms, including those originating from MARL.
 Benchmark scenarios are limited to combat with a limited number of units.
-Those scenarios do not account for the economic part of the game, the so-called “macro” part of the game – collecting resources and building structures.
+Those scenarios do not account for the economic aspect of the game, the so-called “macro” part of the game – collecting resources and building structures.
 Resources in the game (gas and minerals) are required to build units and structures.
-Structures allow for the recruitment of units and their upgrading.
+Structures enable the recruitment of units and their subsequent upgrading.
 The considered scenarios do not involve “worker” units, which collect resources.
-Instead, the scenarios focus on ’micro,’ the term used in the StarCraft II community for unit control during combat.
-The challenge for a human player is multitasking between these two tasks – developing the game economy and controlling the fighting units.
-SMAC focuses only on combat, thereby ensuring this benchmark repeatability, as each iteration involves a fight between the same specification units, eliminating the uncertainty with unit upgrades and different strategies for the economic part of the game.
+Instead, the scenarios focus on ’micro’, a term used in the StarCraft II community to describe unit control during combat.
+In SMACv2, the new addition is randomized maps defined by race (same for both teams), number of units in each team, thus map names like zerg_10_vs_11.
+The new maps randomize types of units according to the probability of each unit type defined in scenario configuration files called capabilities config.
+In addition, the configuration allows for specifying the units’ start position, their spread, and configuration – either surround, reflect, or random probability of each.
+The challenge for a human player is to multitask between these two tasks – developing the game economy and controlling the fighting units.
+SMAC focuses solely on combat, thereby ensuring repeatability of this benchmark, as each iteration involves a fight between the same specification units, eliminating uncertainty due to unit upgrades and different strategies in the economic part of the game.
 In StarCraft II, there are three races, and our experiments covered controlling units of all of them.
-The way to run SMAC is described in our repository.
+The process for running SMAC is described in our repository.
 Appendix B Detailed Results
-Figures 6, 7, and 8 contain more detailed results of our experiments described in Section 4.3.
+Figures 7, 8, and 9 contain more detailed results of our experiments described in Section 4.3.
 The curves on the plots represent the mean percentage test game win rate during the algorithm training for 6 different environments and various mixers.
@@
 For the QPLEX mixer, our method performs much better than MAIC on the 2c_vs_64zg and 27m_vs_30m maps, slightly better on the 5m_vs_6m, slightly worse on MMM2, learns slower but achieves similar final result on the 6h_vs_8z and 3s5z_vs_3s6z maps.
-Our method combined with the VDN mixer gives significantly better results than MAIC on the 3s5z_vs_3s6z map, slightly better on 6h_vs_8z, 27m_vs_30m, 5m_vs_6m, marginally better on the 2c_vs_64zg map and marginally worse on the MMM2 map.
-Figure 6: Means and standard deviations of the percentage wins in the test games during training for QMIX with MAIC, MACTAS, and bare.
-Figure 7: Means and standard deviations of the percentage wins in the test games during training for QPLEX with MAIC, MACTAS, and bare.
-Figure 8: Means and standard deviations of the percentage wins in the test games during training for VDN with MAIC, MACTAS, and bare.
+Our method combined with the VDN mixer gives significantly better results than MAIC on the 3s5z_vs_3s6z map, slightly better on 6h_vs_8z, 27m_vs_30m, 5m_vs_6m, marginally better on the 2c_vs_64zg map, and marginally worse on the MMM2 map.
+Figure 7. Means and standard deviations of the percentage wins in the test games during training for QMIX with MAIC, MACTAS, and bare.
+Figure 8. Means and standard deviations of the percentage wins in the test games during training for QPLEX with MAIC, MACTAS, and bare.
+Figure 9. Means and standard deviations of the percentage wins in the test games during training for VDN with MAIC, MACTAS, and bare.
 Appendix C Tactics used by MACTAS
-The SMAC challenge forces agents to learn various tactics, such as attacking from distance melee units (’kiting’) or using movement advantage (’hit and run’).
+The SMAC challenge forces agents to learn various tactics, such as attacking from a distance with melee units (’kiting’) or using a movement advantage (’hit and run’).
 In this section, we provide an overview of the tactics used by our model based on our understanding of StarCraft II.
 We observed differences in the behavior of our model depending on the random seed.
-The description of tactics is based on replays of our models for seed = 1.
-C.1 2c_vs_64zg
-Figure 9: 2c_vs_64zg replay.
+The description of tactics is based on replays of our models for a seed value of 1.
+C.1. 2c_vs_64zg
+Figure 10. 2c_vs_64zg replay.
 In this scenario, agents control 2 colossi units.
@@
 In this way, the colossi avoid being surrounded and thereby defeated.
-C.2 27m_vs_30m and 5m_vs_6m
-Figure 10: 27m_vs_30m replay.
-Figure 11: 5m_vs_6m replay.
+C.2. 27m_vs_30m and 5m_vs_6m
+Figure 11. 27m_vs_30m replay.
+Figure 12. 5m_vs_6m replay.
 The 27m_vs_30m and the 5m_vs_6m scenarios consist of marine units only.
@@
 In Figure 11, the agent’s units place themselves in a semicircle formation, also known as flanking, which maximizes the number of shooting units, while the enemies block each other.
-Units controlled by MACTAS focus fire on enemy units with lower health, in order to quickly reduce the number of opponents.
+Units controlled by MACTAS focus fire on enemy units with lower health to quickly reduce the number of opponents.
 Furthermore, units with low health retreat and continue fighting, while avoiding being shot.
@@
 However, 27m_vs_30m has a larger number of units that need to cooperate.
-C.3 MMM2
-Figure 12: MMM2 replay.
-The MMM2 is a scenario with a diverse set of units containing 1 medivac (medic), 2 marauders, and 7 marines fighting a team with 1 more marauder and 1 more marine.
+C.3. MMM2
+Figure 13. MMM2 replay.
+The MMM2 is a scenario with a diverse set of units containing 1 medivac (medic), 2 marauders, and 7 marines fighting against a team with 1 more marauder and 1 more marine.
 Medivac cannot attack, but can heal units as long as it has energy.
@@
 In this scenario, we can observe similar behavior in the MACTAS medivac control, which results in winning combat against an opponent who has a numerical advantage.
-C.4 6h_vs_8z
-Figure 13: 6h_vs_8z replay.
+C.4. 6h_vs_8z
+Figure 14. 6h_vs_8z replay.
 Agents control 6 range shooting units, hydralisks, while the enemy controls 8 melee zealots.
@@
 Notably, the zealot units within this scenario are not upgraded with the “charge” ability, which would otherwise allow them to have a significant increase in movement speed and easily reach hydralisks.
-C.5 3s5z_vs_3s6z
-Figure 14: 3s5z_vs_3s6z replay.
+C.5. 3s5z_vs_3s6z
+Figure 15. 3s5z_vs_3s6z replay.
 MACTAS controls 3 stalkers and 5 zealots, while the enemy has one more zealot.
 The stalkers are fast, range-attacking units, but with weak defense.
-On the other hand, zealots have good defense, but need to get close to the opponent unit.
+On the other hand, zealots have a good defense, but they need to get close to the opponent’s unit.
 The tactic in this scenario is interesting.
@@
 Appendix D Sensitivity analysis
-Figure 15: MACTAS; sensitivity of the results to architecture size with the QMIX mixer.
-Figure 16: MACTAS; sensitivity of the results to architecture size with the QPLEX mixer.
-Figure 17: MACTAS; sensitivity of the results to architecture size with the VDN mixer.
-Due to the technical issues, the version with 3 Transformer layers and a width of the Transformer feedforward layer equal to 512 was evaluated on 4 seeds of the random generator (1 to 4).
-Figure 18: MACTAS; sensitivity of the results to the dropout parameter in the Transformer block. The default value of this parameter is $10\%$.
-Figure 19: MACTAS; sensitivity of the top 2 exploration method to the temperature coefficient.
-We evaluated the choice of the number of Transformer encoder layers and the width of the Transformer feedforward layer for MACTAS with QMIX, QPLEX, and VDN as a mixer in five environments: 2c_vs_64zg, 5m_vs_6m, 27m_vs_30m, MMM2 and 6h_vs_8z.
+Figure 16. MACTAS; sensitivity of the results to architecture size with the QMIX mixer.
+Figure 17. MACTAS; sensitivity of the results to architecture size with the QPLEX mixer.
+Figure 18. MACTAS; sensitivity of the results to architecture size with the VDN mixer.
+Figure 19. MACTAS; sensitivity of the results to the dropout parameter in the Transformer block. The default value of this parameter is $10\%$.
+Figure 20. MACTAS; sensitivity of the top 2 exploration methods to the temperature coefficient.
+We evaluated the choice of the number of Transformer encoder layers and the width of the Transformer feedforward layer for MACTAS with QMIX, QPLEX, and VDN as a mixer on six SMAC(v1) maps: 2c_vs_64zg, 5m_vs_6m, 27m_vs_30m, MMM2, 6h_vs_8z, and 3s5z_vs_3s6z.
 We experimented with the dropout of the Transformer encoder on all six maps.
-We also assessed the sensitivity of our new top $k$ exploration technique to the softmax temperature in the case of MACTAS+QMIX and $k=2$ on three maps: 27m_vs_30m, MMM2 and 3s5z_vs_3s6z.
+We also assessed the sensitivity of our new top $k$ exploration technique to the softmax temperature in the case of MACTAS+QMIX and $k=2$ on all six maps.
 […]
@@
 We observe that this method is sensitive to this parameter.
-For the 27m_vs_30m and MMM2 maps, the low temperatures work better, with $0.25$ achieving similar results to the $\epsilon$-greedy scheme, which is equivalent to choosing 0 as the temperature.
-In these cases, the new exploration scheme does not outperform the $\epsilon$-greedy exploration.
+For the 2c_vs_64zg, 5m_vs_6m, 27m_vs_30m, MMM2, and 6h_vs_8z maps, the low temperatures ($0.25$, $0.33$) work best with better results for the $0.25$ temperature.
 We can observe an interesting situation on the 3s5z_vs_3s6z map.
-The top 2 exploration scheme outperforms the $\epsilon$-greedy exploration for the $0.33$ temperature.
-However, a temperature equal to $4$ gives a similar result to the $\epsilon$-greedy exploration.
-The in-between temperatures perform worse than both of these values, showing that there is no monotonic relationship between the temperature and the performance.
+The best results are obtained for the $0.33$ temperature, with the temperature $4$ being the second-best, and the temperature $0.25$ performing quite poorly, which suggests that, in general, the relationship between the temperature value and performance may not be monotonic.
+For the most challenging maps, which require 7.5 million training steps, the top 2 exploration schemes outperform the $\epsilon$-greedy exploration, which is equivalent to choosing 0 as the temperature.
+In other cases, the new exploration scheme does not outperform the $\epsilon$-greedy exploration; however, except for the case of the 5m_vs_6m map, where training is slowed down, the method remains indifferent.
 Appendix E Software
 Table 3 contains a list of key programming libraries used during our experiments.
-The complete list of the programming libraries is available in the MACTAS’ official code repository in file requirements.txt.
+The complete list of programming libraries is available in the MACTAS’ official code repository in the file requirements.txt.
 Library name Version
@@
 torch 2.5.1
-Table 3: Essential libraries in programming environment.
+Table 3. Essential libraries in programming environment.
 Appendix F Hardware
@@
 Cluster 3 NVIDIA A100 PCIe 80GB 565.57.01 AMD EPYC 7713 64-Core Processor Linux-6.8.0-64-generic-x86_64-with-glibc2.39 2048 GB
-Table 4: Specification of the computational clusters used in the experiments.
-Here we present the hardware used in our experiments. We used NVIDIA Ampere GPU architecture on three computational clusters. The specific setup is presented in Table 4.
+Table 4. Specification of the computational clusters used in the experiments.
+Here we present the hardware used in our experiments. We utilized the NVIDIA Ampere GPU architecture across three computational clusters. The specific setup is presented in Table 4.

### v3

@@
-MACTAS: Self-Attention-Based Module for Inter-Agent Communication in Multi-Agent Reinforcement Learning
-Abstract.
+MACTAS: Self-Attention-Based Inter-Agent Communication in Multi-Agent Reinforcement Learning with Action-Value Function Decomposition
+Abstract
 Communication is essential for the collective execution of complex tasks by human agents, motivating interest in communication mechanisms for multi-agent reinforcement learning (MARL).
 However, existing communication protocols in MARL are often complex and non-differentiable.
-In this work, we introduce a self-attention-based communication module that exchanges information between the agents in MARL.
+In this work, we introduce a self-attention-based communication method that exchanges information between the agents in MARL.
 Our proposed approach is fully differentiable, allowing agents to learn to generate messages in a reward-driven manner.
-The module can be seamlessly integrated with any action-value function decomposition method and can be viewed as an extension of such decompositions.
-Notably, it includes a fixed number of trainable parameters, independent of the number of agents.
-Experimental results on the SMAC and SMACv2 benchmarks demonstrate the effectiveness of our approach, which achieves state-of-the-art performance on a number of maps.
-Key words and phrases:
-Multi-Agent Reinforcement Learning, Cooperation in Multi-Agent Reinforcement Learning, Communication in Multi-Agent Reinforcement Learning
-1. Introduction
+The method can be seamlessly integrated with any action-value function decomposition algorithm and can be viewed as an orthogonal extension of such decompositions.
+Notably, it includes a fixed number of trainable parameters, independent of the number of agents, which makes it scalable to large systems.
+Experimental results on the SMACv2 benchmark demonstrate the effectiveness of our approach, which achieves state-of-the-art performance on a number of maps.
+**footnotetext: Equal contribution.
+1 Introduction
 Multi-agent reinforcement learning (MARL) Ning and Xie (2024); Du and Ding (2021); Nguyen et al.
 (2020); Oroojlooy and Hajinezhad (2023); Wong et al.
-(2023) promises efficient control in large-scale systems through the collective learning of many agents in dynamic environments.
-Multiple agents interact within a shared environment to learn optimal policies.
-MARL is defined as the process in which agents optimize their behaviors based on the history of their interactions with the environment, i.e., on their experience.
-Research in this area aims to design methods that enable agents to learn quickly, i.e., from minimal experience.
-The challenges and complexities in using MARL for controlling multi-agent systems include coordination among agents, non-stationary environments due to evolving agent policies, and scalability issues.
-Advanced MARL approaches often involve sophisticated techniques, including decentralized training, communication protocols among agents, and hierarchical methods to manage multi-agent interactions as effectively as possible.
-Theoretical advancements and practical applications of MARL demonstrate its potential in diverse fields ranging from robotics Orr and Dutta (2023) and autonomous driving Yadav et al.
+(2023) offers a powerful framework for controlling large-scale systems by enabling multiple agents to learn collectively while interacting within a shared, dynamic environment.
+In MARL, agents seek to optimize their policies based on accumulated experience, derived from the history of their interactions with both the environment and other agents.
+A central objective of MARL research is to develop algorithms that achieve effective learning with minimal experience, thereby improving sample efficiency.
+Despite its promise, MARL presents several fundamental challenges, including the need for coordination among agents, non-stationarity induced by concurrently evolving policies, and scalability to systems with many agents.
+To address these challenges, advanced MARL methods incorporate a range of techniques such as decentralized or partially centralized training, explicit communication mechanisms, and hierarchical structures that facilitate efficient multi-agent coordination.
+Ongoing theoretical developments and practical successes highlight the versatility of MARL across a broad spectrum of applications, including robotics Orr and Dutta (2023), autonomous driving Yadav et al.
 (2023); Schmidt et al.
 (2022); Althamary et al.
-(2019) to telecommunications Li et al.
+(2019), and telecommunications Li et al.
 (2022b); Feriani and Hossain (2021).
-Figure 1. 3s5z_vs_3s6z scenario in the StarCraft Multi-Agent Challenge (SMAC).
-The most prominent research topics in MARL are the action-value function decompositions Wang et al.
+Figure 1: The strategic sacrifice tactics found by MACTAS for the 3s5z_vs_3s6z map from the SMAC benchmark, on which MACTAS outperforms the baselines significantly.
+The agent at the top of the figure (“Local Player”) lures the opponents away and sacrifices its own life to enable the remaining agents to gain a strategic advantage and win the episode.
+Two of the most actively studied directions in MARL are action–value function decomposition methods Wang et al.
 (2021); Liu et al.
-(2023b), where agents share a joint reward according to their contributions, and inter-agent communication Liu et al.
+(2023b), which allocate a shared team reward according to individual agent contributions, and inter-agent communication mechanisms Liu et al.
 (2023a); Guo et al.
-(2023), where agents exchange information in order to coordinate their joint policy.
-The most challenging multi-agent environments are partially observable, typically by limiting agents’ information to their close proximity, and the achievement of their common goals is predicated upon coordination, which needs to involve some form of communication.
-In most research, the action-value function and inter-agent communication are issues addressed separately.
-Therefore, in this paper, we introduce communication into the very core of the action-value function decomposition.
-The prevailing paradigm in cooperative MARL is Centralized Training and Decentralized Execution (CTDE), where the training algorithm utilizes the global state, but agents can only use their own local observations.
-We argue that in most challenging MARL tasks, this is insufficient.
-Moreover, in most challenging MARL tasks, some form of inter-agent communication is a necessity, which can be provided relatively easily (e.g., through WiFi).
-Therefore, we introduce a communication protocol that collects information from each agent in a central entity and broadcasts the combined information to all agents.
-Because information aggregation is based on attention, our proposed architecture requires a constant number of weights, plus an embedding, for each agent.
-It is easy to apply using readily available and efficient implementations of the Transformer encoder layer.
-The contribution of this paper may be summarized as follows:
-  • We introduce Multi-Agent Communication via Transformer on Agents’ States (MACTAS) module for inter-agent communication in MARL. This module allows the agent to share their state of consciousness.
-  • We verify MACTAS experimentally on several challenging maps from SMACSamvelyan et al.
-(2019) and SMACv2 Ellis et al.
-(2023), giving excellent results and establishing state-of-the-art in some cases.
-2. Related Work
-Single-Agent Reinforcement Learning (SARL) Sutton and Barto (2018) addresses trial-and-error learning in sequential decision-making under uncertainty.
-In Multi-Agent Reinforcement Learning (MARL) Ning and Xie (2024), a group of agents learns to act in the same dynamic environment, collectively modifying its state.
-There are various settings for MARL, depending on whether the agents operate synchronously or asynchronously, whether they have common or adverse goals, and other circumstances.
-In this paper, we focus on the synchronous and cooperative MARL, with a single reward value for all agents at every time instance.
-MARL has been addressed by extending well-established SARL algorithms, including Multi-Agent SAC Wei et al. (2018), Multi-Agent PPO Yu et al. (2022), and Multi-Agent DDPG Lowe et al. (2017).
-Methods of cooperative MARL often adopt a centralized training with decentralized execution (CTDE) paradigm, in which agents share information during training but act based solely on their own observations during execution.
-However, this paradigm suffers from the global action-value function, whose complexity grows exponentially with the number of agents Kraemer and Banerjee (2016).
-Therefore, action-value function decomposition is a fundamental problem in MARL.
-Existing decomposed architectures include modules assigned to agents.
-These modules are fed with agents’ observations and their previous actions and produce agents’ contributions to the architecture’s output.
-The best action for the agent is the one that makes the highest contribution.
-In VDN Sunehag et al.
-(2018), these contributions are just added.
-In QMIX Rashid et al.
-(2018), a neural module, the mixing network, produces its weighted sum with positive weights.
-In Weighted QMIX Rashid et al.
-(2020), the architecture is trained with a special weight assigned to the collective actions deemed the best.
-In QTRAN Son et al.
-(2019), an additional module fed with all the actions of all agents corrects the weighted sum of the agents’ contributions.
-In Qatten Yang et al.
-(2020), the mixing network is based on attention.
-In CDS Li et al.
-(2021), a component of the mixing network optimizes the amount of exploration undertaken by the agents.
-In DVD Li et al.
-(2022a), the internal state of the modules assigned to agents is also passed to the mixing networks.
-In SHAQ Wang et al.
-(2022), the mixing network is enriched by several components based on Shapley value theory.
-In QPLEX Wang et al.
-(2021), a carefully engineered duplex dueling of pairs of contributions of agents is introduced into the mixing network.
-In NA2Q Liu et al.
-(2023b), a similar duplex dueling is performed with the use of attention.
-Another fundamental problem in cooperative MARL is the coordination of agents’ activity for which a certain communication protocol is required.
+(2023), which enable coordination through information exchange.
+Many of the most challenging multi-agent environments are only partially observable, often restricting agents’ perceptions to their local neighborhoods.
+In such settings, successful task completion critically depends on effective coordination, which in turn typically requires some form of communication.
+In our proposed architecture, communication is integrated directly into the action–value function decomposition itself.
+The dominant framework for cooperative MARL is Centralized Training and Decentralized Execution (CTDE), wherein the learning algorithm has access to the global state during training, while each agent relies solely on its local observations at execution time.
+We contend that this paradigm is inadequate for many complex MARL problems.
+In practice, reliable inter-agent communication is often feasible—for example, via wireless networks—and can significantly enhance coordination.
+Motivated by this observation, we introduce a communication mechanism based on self-attention: The latent state of each agent is updated based on other agents’ states, similar to how word embeddings are updated based on other words in the Transformer encoder Vaswani et al.
+(2017).
+By leveraging self-attention, the proposed architecture requires only a constant number of shared parameters, along with agent-specific embeddings.
+Furthermore, it can be efficiently implemented using standard Transformer encoder layers, making it both scalable and practical.
+Although self-attention has been employed in different contexts in MARL before, including communication, this is the first successful approach to using self-attention on its own as a communication tool in MARL, without the need for graph-structure learning or global state reconstruction.
+Moreover, our method can be combined with an arbitrary action-value decomposition.
+None of the previous works has applied self-attention directly as a communication protocol.
+Our work fills this gap.
+It is the first algorithm showing that self-attention on agent internal states is sufficient for effective and reliable communication without the need for explicit teammate modeling or message preparation mechanisms.
+The contribution of this paper includes:
+  • We introduce Multi-Agent Communication via Transformer on Agents’ States (MACTAS) for inter-agent communication in MARL based on a self-attention mechanism applied directly to internal state embeddings of agents.
+MACTAS is the first self-attention-based communication mechanism with the number of trainable parameters independent of the number of agents, which makes it scalable to large systems.
+Moreover, it can be combined with any action-value function mixing architecture.
+  • We verify experimentally MACTAS in combination with the most popular decomposition methods on challenging maps from SMACv2, showing results better than current state-of-the-art algorithms in most cases.
+We also perform experiments on SMAC, Level-based Foraging, and Hallway, proving that MACTAS is more reliable in a diverse set of scenarios.
+Our approach outperforms state-of-the-art communication algorithms (including a method that uses self-attention and global state reconstruction) in combination with common mixers.
+Moreover, our method proves robust on a diverse set of tasks, while each baseline fails on some relatively simple task (like Hallway or the 2c_vs_64zg SMAC map).
+Our algorithm can learn sophisticated tactics, such as sacrificing one agent to win the entire battle (as shown in Figure 1).
+A detailed study of the tactics found by our method is presented in Appendix E.
+2 Related Work
+Single-agent reinforcement learning (SARL) Sutton and Barto (2018) studies trial-and-error learning for sequential decision-making under uncertainty.
+Multi-agent reinforcement learning (MARL) Ning and Xie (2024) generalizes this framework to settings in which multiple agents simultaneously interact with and influence a shared, dynamic environment.
+MARL encompasses a wide range of problem formulations, varying along dimensions such as synchronous versus asynchronous decision-making, cooperative versus competitive objectives, and differing information structures.
+In this work, we restrict our attention to the synchronous, fully cooperative setting, where all agents share a common reward signal at each time step.
+A major limitation of the CTDE framework is the exponential growth of the joint action–value function with the number of agents Kraemer and Banerjee (2016).
+As a result, action–value function decomposition has emerged as a central challenge in MARL.
+Decomposition-based architectures typically assign each agent a local module that processes its observations and previous actions to produce an individual contribution to the global value.
+An agent’s optimal action is then determined by maximizing its contribution to the joint value estimate.
+Several decomposition approaches have been proposed in the literature.
+Value Decomposition Networks (VDN) Sunehag et al.
+(2018) aggregate agent contributions through simple summation, while QMIX Rashid et al.
+(2018) employs a mixing network that computes a weighted sum with non-negative weights to preserve monotonicity.
+Weighted QMIX Rashid et al.
+(2020) further refines this approach by assigning special weights to optimal joint actions during training.
+QTRAN Son et al.
+(2019) introduces an auxiliary module that incorporates the joint action to correct the decomposed value.
+Attention mechanisms are leveraged in Qatten Yang et al.
+(2020) to construct a more flexible mixing network.
+Additional extensions include CDS Li et al.
+(2021), which adjusts agent exploration through the mixing network; DVD Li et al.
+(2022a), which incorporates agents’ internal states into value aggregation; SHAQ Wang et al.
+(2022), which augments the mixing network using principles from Shapley value theory; QPLEX Wang et al.
+(2021), which introduces a duplex dueling architecture over agent contributions; and NA2Q Liu et al.
+(2023b), which adopts a similar duplex dueling strategy enhanced with attention mechanisms.
+Another fundamental problem in cooperative MARL is the coordination of agents’ activities, which requires a communication protocol.
 A number of solutions have been proposed for this purpose: CommNet Sukhbaatar et al.
@@
 (2020) uses two-stage attention mechanism to limit number of messages exchanged between agents, IS Kim et al.
-(2021) encodes trajectory of future actions for each agent and shares them with others, in MAIC Yuan et al.
-(2022) agents explicitly model their teammates and use it to bias their value function, DHCG Liu et al.
+(2021) encodes trajectory of future actions for each agent and shares them with others, DHCG Liu et al.
 (2023a) learns a dependency graph between agents and uses it to limit the number of communication edges, TEM Guo et al.
-(2023) uses a Transformer to simulate communication resembling email exchange, where agents only directly communicate with teammates nearby, but can forward messages from one to another, ACE Li et al.
-(2023) proposes bidirectional action-dependent Q-learning.
+(2023) uses a Transformer to simulate communication resembling email exchange, where agents only directly communicate with teammates nearby, but can forward messages from one to another, ExpoComm Li et al.
+(2025) utilizes exponential graphs to design the communication topology.
+Also, in MAIC Yuan et al.
+(2022) agents explicitly model their teammates and use it to bias their value function, MASIA Guan et al.
+(2022) uses attention-based communication architecture with a decoder module to reconstruct the global state, CommFormer Hu et al.
+(2024) uses the attention encoder and decoder architecture combined with learning a communication graph, having the number of parameters independent of the number of agents.
 In addition, there are solutions in which the agents communicate in natural language, including: TWOSOME Tan et al.
@@
 (2024) which separates the output of the LLM into communication and actions, allowing LLM agents to exchange interpretable verbal messages.
-We acknowledge various applications of attention or Transformer architectures in MARL algorithms, such as in Qatten mixing network, or in communication algorithms like NA2Q, GA-Comm, and TEM.
-None of the previous works has applied the Transformer architecture directly as a communication protocol.
-Our work fills this gap.
+Attention and Transformer architectures have been applied in MARL algorithms, such as in Qatten and NA2Q mixing networks, or in communication mechanisms like GA-Comm, TEM, CommFormer, and MASIA.
+However, our approach is conceptually different: We apply the attention mechanism directly to generate the agents’ messages, without additional graph-structure learning (as in CommFormer) or decoding the global state (as in MASIA).
+Our proposed use of attention makes the number of trainable parameters independent of the number of agents (which enables deployment at scale), unlike in MASIA.
+Moreover, our method can be combined with any mixing architecture, unlike CommFormer.
 We consider communication methods that can be coupled with an arbitrary action-value decomposition method.
 We also allow agents to send messages based only on their local information (current observation, historical observations, and historical actions).
-In this framework, the best results so far have been achieved by MAIC Yuan et al.
+In this framework, MAIC has achieved the best results so far.
+MASIA is a communication protocol that uses self-attention.
+Therefore, in the experimental study, we compare our proposed MACTAS to both MAIC and MASIA.
+3 Method
+In this section, we formalize the problem statement, present the details of applying self-attention in MACTAS, and discuss its distributed deployment.
+Figure 2: The proposed MACTAS architecture.
+Problem definition.
+In this paper, we discuss the problem of MARL using the formalism of a cooperative Decentralized Partially Observable Markov Decision Process (Dec-POMDP) with communication.
+Dec-POMDP is defined by a tuple $\langle\mathbb{N},\mathbb{S},\mathbb{A},\mathcal{R},\mathcal{P},\Omega,\mathcal{O},\gamma\rangle$, where $\mathbb{N}=\{1,\dots,n\}$ represents the team of agents, $\mathbb{S}$ is the global environmental state space, $\mathbb{A}=\mathbb{A}^{1}\times\dots\times\mathbb{A}^{n}$ is the joint action space, $\mathcal{R}:\mathbb{S}\times\mathbb{A}\times\mathbb{S}\rightarrow\mathbb{R}$ is the reward function, $\mathcal{P}:\mathbb{S}\times\mathbb{A}\rightarrow P(\mathbb{S})$ is the state transition probability function, $\Omega=\Omega^{1}\times\dots\times\Omega^{n}$ is the joint observation space, $\mathcal{O}:\mathbb{N}\times\mathbb{S}\rightarrow\Omega$ is the observation function and $\gamma\in[0,1]$ is the discount factor.
+At time $t=1,2,\dots$ the $i$-th agent makes an observation, $o^{i}_{t}=\mathcal{O}(i,s_{t}),o^{i}_{t}\in\Omega^{i}$, of the environment state $s_{t}\in\mathbb{S}$.
+Agents may exchange messages through a channel that generally refines them with some processing.
+We define a set of representations $\mathcal{J}=\mathcal{J}^{1}\times\dots\times\mathcal{J}^{n}$.
+Each agent at timestep $t$ maintains a representation, $\mathcal{I}^{i}_{t}\in\mathcal{J}^{i}$, of observations it has made, messages it has received, and actions it has taken.
+Based on thereof and a policy, $\pi^{i}:\mathcal{J}^{i}\rightarrow\mathbb{A}^{i}$, the agent performs an action, $a^{i}_{t}\in\mathbb{A}^{i}$, $a^{i}_{t}\sim\pi^{i}(\cdot|\mathcal{I}^{i}_{t})$.
+The team action $a_{t}=[(a^{1}_{t}),\dots,(a^{n}_{t})]$ impacts the next environment state $s_{t+1}\sim\mathcal{P}(\cdot|s_{t},a_{t})$ and the reward $r_{t}=\mathcal{R}(s_{t},a_{t},s_{t+1})$ is given collectively to all agents.
+The goal of the agents’ learning is to optimize their team policy $\pi=[\pi^{1},\dots,\pi^{n}]$ to expect in each environment state $s$ the highest sum of discounted rewards:
+$\mathcal{V}^{\pi}(s)=\mathbb{E}\left(\sum_{i\geq 0}\gamma^{i}r_{t+i}\Big|s_{t}=s,\text{policy in use}=\pi\right).$
+The goal of this paper is to establish the message exchange protocol that supports the above learning goal.
+Self-attention as communication protocol.
+Suppose at time step $t$, each agent $i$ maintains a latent state $h_{t}^{i}\in\mathbb{R}^{n_{h}}$ (we use the horizontal vector notation) derived from its observations up to that point.
+To select an optimal action, an agent should incorporate information about other agents’ states.
+To facilitate such information sharing, we employ a self-attention mechanism Vaswani et al.
 (2017).
-3. Method
-Figure 2. The proposed MACTAS architecture.
-3.1. Problem definition
-In this paper, we discuss the problem of MARL using the formalism of a cooperative Decentralized Partially Observable Markov Decision Process (Dec-POMDP) with communication.
-Dec-POMDP is defined by a tuple $\langle\mathbb{N},\mathbb{S},\mathbb{A},\mathcal{R},\mathcal{P},\Omega,\mathcal{O},\gamma\rangle$, where $\mathbb{N}=\{1,\dots,n\}$ represents the team of agents, $\mathbb{S}$ is the global environmental state space, $\mathbb{A}=\mathbb{A}^{1}\times\dots\times\mathbb{A}^{n}$ is the joint action space, $\mathcal{R}$ is the reward function, $\mathcal{P}$ is the state transition probability function, $\Omega=\Omega^{1}\times\dots\times\Omega^{n}$ is the joint observation space, $\mathcal{O}$ is the observation function and $\gamma$ is the discount factor.
-At time $t=1,2,\dots$ each $i$-th agent makes an observation, $o^{i}_{t}=\mathcal{O}^{i}(s_{t}),o^{i}_{t}\in\Omega^{i}$, of the environment state $s_{t}\in\mathbb{S}$.
-Agents may exchange messages with others through a certain channel that may generally refine these messages with some processing.
-Each agent maintains a representation, $\mathcal{I}^{i}_{t}$, of observations it has made, messages it has received, and actions it has taken.
-Based on thereof and a policy, $\pi^{i}$, the agent performs an action, $a^{i}_{t}\in\mathbb{A}^{i}$, $a^{i}_{t}\sim\pi^{i}(\cdot|\mathcal{I}^{i}_{t})$.
-The team action $a_{t}=[(a^{1}_{t}),\dots,(a^{n}_{t})]$ impacts the next environment state $s_{t+1}\sim\mathcal{P}(\cdot|s_{t},a_{t})$ and the reward $r_{t}=\mathcal{R}(s_{t},a_{t},s_{t+1})$ is given collectively to all agents.
-The goal of the agents’ learning is to optimize their team policy $\pi=[\pi^{1},\dots,\pi^{n}]$ to expect in each environment state $s$ the highest sum of discounted rewards
-$\mathcal{V}^{\pi}(s)=\mathbb{E}\left(\sum_{i\geq 0}\gamma^{i}r_{t+i}\Big|s_{t}=s,\text{policy in use}=\pi\right),$
-where $\gamma\in[0,1]$.
-The goal of this paper is to establish the message exchange protocol that supports the above goal of learning.
-3.2. MACTAS
+Specifically, let $M^{k},M^{q},M^{v}\in\mathbb{R}^{n_{h}\times n_{h}}$ denote learnable projection matrices.
+Agent $i$ updates its representation by aggregating the states of all agents, where the contribution of agent $j$ is given by $\alpha^{i,j}_{t}h^{j}_{t}M^{v}$.
+The attention weight $\alpha^{i,j}_{t}$ is proportional to $\exp\left(h^{i}_{t}M^{q}(h^{j}_{t}M^{k})^{T}/\sqrt{n^{h}}\right)$, and is normalized such that $\sum_{j}\alpha^{i,j}_{t}=1$.
+This weighting reflects the relevance of agent $j$’s state to agent $i$ at time $t$.
+The Transformer encoder introduced in Vaswani et al.
+(2017) augments self-attention with several additional components.
+These include (i) multi-head attention, which allows different subspaces of the state representations to be attended to in parallel; (ii) a feedforward MLP applied to the attention outputs; (iii) layer normalization applied after residual updates; and (iv) stacking multiple such blocks to form deep architectures.
+We adopt this Transformer encoder structure to enable effective and flexible information exchange among agents.
+MACTAS.
 Our proposed Multi-Agent Communication via Transformer on Agents States (MACTAS) architecture is presented in Fig.
@@
 The input module $f^{i}$ usually comprises a feedforward part, denoted by MLP in Figure 2, a recurrent part, denoted by GRU, and the output feedforward part, again denoted by MLP.
-Intuitively, we understand the $f^{i}$ module as the consciousness of the $i$-th agent.
+Intuitively, we understand the $f^{i}$ module as the “awareness” of the $i$-th agent.
 Based on what the agent is aware of, it determines its contribution to future discounted rewards, depending on the action it currently takes.
-In order to give a global view on the current state to the agent’s consciousness, we introduce a communication mechanism.
-We divide the $f^{i}$ module after its recurrent part and feed its state $h^{i}_{t}\in\real^{n_{h}}$ to the module $g$, where the states of consciousness of the agents are confronted.
-As a result, the $g$ module produces increments, $z^{i}_{t}\in\real^{n_{h}}$.
+In order to give a global view on the current state to the agent’s “awareness”, we introduce a communication mechanism.
+We divide the $f^{i}$ module after its recurrent part and feed its state $h^{i}_{t}\in\mathbb{R}^{n_{h}}$ to the module $g$, where the states of “awareness” of the agents are confronted.
+As a result, the $g$ module produces increments, $z^{i}_{t}\in\mathbb{R}^{n_{h}}$.
 The sum $z^{i}_{t}+h^{i}_{t}$ is then passed to the second MLP part of the $f^{i}$ module.
-Consequently, the consciousness of each agent is based not only on its observations and previous actions, but also on observations and previous actions of all other agents.
+Consequently, the “awareness” of each agent is based not only on its observations and previous actions, but also on observations and previous actions of all other agents.
 The minimal requirement for the $g$ module is as follows:
-  1. (1)
-    It is fed with $n$ vectors and produces $n$ vectors, both of size $\dim(h^{i}_{t})$. $i$-th input vector corresponds to the $i$-th output vector.
-  2. (2)
+  1. 1.
+    It is fed with $n$ vectors and produces $n$ vectors, both of size $\dim(h^{i}_{t})$. The $i$-th input vector corresponds to the $i$-th output vector.
+  2. 2.
     It is a layered structure and the layers include self-attention Vaswani et al. (2017).
@@
 (2017), with the following settings:
-  1. (1)
+  1. 1.
     There are no positional embeddings. The agents’ embeddings are fed to $f^{i}$ instead.
-  2. (2)
+  2. 2.
     There is no masking.
-  3. (3)
-    We use a separate Adam optimizer Kingma and Ba (2017) for the Transformer parameters; the rest of the architecture is trained with its originally proposed optimizer (typically it is RMSProp Tieleman and Hinton (2012)).
-  4. (4)
-    The weights of the output layer of the top Transformer module are initialized with zeros. Therefore, initially, this module does not impact the state of consciousness of the agents.
-A good performance of the Transformer in this role is an expected result.
+  3. 3.
+    We use a separate Adam optimizer Kingma and Ba (2017) for the Transformer parameters; the rest of the architecture is trained with its originally proposed optimizer (typically RMSProp Tieleman and Hinton (2012)).
+  4. 4.
+    The weights of the output layer of the top Transformer module are initialized with zeros. Therefore, initially, this module does not affect the agents’ state of “awareness”.
+A good performance of the Transformer in this role is expected.
 This architecture has been specifically optimized to transform a set of embeddings into another set of the same size, which is exactly what we require in the $g$ module.
-3.3. Implementation considerations
+Distributed deployment.
 The basic approach to implement a multi-agent system with our proposed MACTAS module is to set up the Intermediate Processing Unit (IPU) on which the module is run.
@@
 Then, each agent needs to take on the part of processing that consists of refining $h^{i}_{t}$ and results in $z^{i}_{t}$.
-That requires exchange of messages with other agents as many times as many Transformer layers are stacked in MACTAS, each time with $O(n)$ messages; $O(n^{2})$ for all agents.
+That requires exchanging messages with other agents as many times as there are Transformer layers in MACTAS, each time with $O(n)$ messages; $O(n^{2})$ for all agents.
 In training, the agents need to aggregate increments in their shared weights.
-In the distributed implementation, it may happen that some of the agents are out of reach of others and do not exchange information with them.
+In a distributed implementation, some agents may be out of reach of others and not exchange information with them.
 That only means that self-attention in MACTAS is forced to assign zero weight to unavailable transmissions from those unheard agents.
-3.4. Exploration
-Learning in a multi-agent system with decentralized execution is hampered by agents’ policies falling into suboptimal Nash equilibria.
-Let us consider a typical MARL setting where agents use the $\epsilon$-greedy strategy with $\epsilon$ annealing and explore independently.
-On average, only $n\epsilon$ agents explore.
-To learn that a certain combination of agents’ actions is better than the current Nash equilibrium, at least two agents need to explore and try a better combination.
-However, this happens rarely.
-What happens more often is that only one agent explores, which leads to a combination worse than greedy.
-Note that this effect may be even stronger for a better approximation of the $Q$ function, as it leaves less room for accidental switches of the greedy action combinations.
-To counteract this effect, we propose a combination of $\epsilon$-greedy and Boltzmann exploration.
-Namely, each agent with probability $\epsilon$ chooses independently a random action; however, with probability $1-\epsilon$ it does not act greedily, but chooses from the top $k$ of its actions (according to its local action-value function) using the Boltzmann strategy.
-Using $k=1$ is equivalent to the pure $\epsilon$-greedy policy.
-Selecting a small $k$ leads to exploring a region that is quite good according to the agent.
-We observe that $k=2$ works well in practice.
-The temperature of the Boltzmann exploration controls the amount of extra exploration that we want to add.
-Hyperparameter Value
-Batch size 32
-Test episodes 32
-Replay buffer size 5000
-Discount factor 0.99
-Start epsilon 1.0
-Finish epsilon 0.05
-Anneal steps 50,0001 or 100,0002
-RNN units 64
-Optimizer RMSProp
-Learning rate 0.0005
-Table 1. Basic experimental settings and hyperparameters. 1 for 2.5 mln. step maps. 2 for 7.5 mln. step maps.
 Scenario Stacked encoders FFN dimension
+all SMACv2 maps 3 512
 2c_vs_64zg 2 512
-5m_vs_6m 1 128
-27m_vs_30m 1 128
+5m_vs_6m, 27m_vs_30m 1 128
 MMM2 3 256
-6h_vs_8z 4 512
-3s5z_vs_3s6z 3 512
-All SMACv2 scenarios 3 512
-Table 2. MACTAS shape for each scenario.
-Figure 3. Percentage of wins in test games in training time for the connection protocols MACTAS and MAIC, and the $Q$ architectures QMIX, QPLEX, and VDN.
-Figure 4. Percentage of wins and standard deviations in test games in training time for the connection protocols MACTAS, MAIC, and the bare mixer for the $Q$ architecture QMIX
-Figure 5. The impact of our exploration scheme on MACTAS+QMIX.
-Figure 6. The impact of our residual connection on MACTAS+QMIX.
-4. Experiments
-In this section, we present the results of our proposed MACTAS communication protocol in the StarCraft Multi-Agent Challenge environments (SMAC Samvelyan et al.
-(2019) and SMACv2 Ellis et al.
+6h_vs_8z, 3s5z_vs_3s6z 3 512
+LBF and Hallway 1 32
+Table 1: MACTAS shape for each scenario.
+4 Experiments
+Figure 3: Training curves for the zerg race on the SMACv2 benchmark with the QMIX mixer.
+In this section, we present the results of our proposed MACTAS communication protocol in the StarCraft Multi-Agent Challenge environments (SMACv2 Ellis et al.
+(2023) and SMAC Samvelyan et al.
 (2019)).
 We compare it with the state-of-the-art MAIC Yuan et al.
-(2022), in combinations with 3 mixer algorithms: QMIX, QPLEX, and VDN, on 6 maps.
-We present results of two ablation studies regarding the design of neural network connectivity and the exploration form.
-Additionally, in the Appendix, we report the sensitivity study of MACTAS to hyperparameters, lower-level SMAC and SMACv2 settings, and analyses of the most interesting strategies that have emerged in training.
-4.1. StarCraft Multi-Agent Challenge
+(2022) and MASIA Guan et al.
+(2022), in combination with 3 mixer algorithms: QMIX, QPLEX, and VDN, on 9 SMACv2 maps and in combination with QMIX on 6 SMAC maps.
+We also present results in combination with QMIX on Level-based Foraging Papoudakis et al.
+(2021) and Hallway Wang et al.
+(2019).
+4.1 StarCraft Multi-Agent Challenge
 The StarCraft Multi-Agent Challenge (SMAC and SMACv2) is a micromanagement benchmark where two opposing teams face each other in different scenarios (maps), with one team controlled by the RL agent and the other by a built-in algorithm.
 Maps differ by difficulty, mainly due to an imbalance in the number of controlled units.
-For our experiments, we selected 2 hard and 4 super-hard SMAC scenarios, which are widely benchmarked in the MARL community:
-  • 2c_vs_64zg with 2 colossus units leveraging terrain advantage against 64 small enemy zerglings. Level: hard.
-  • 5m_vs_6m with 5 cooperating marines and 6 enemy marines. Level: hard.
-  • 27m_vs_30m big battle of 27 marines conquering 30 marines. Level: super-hard.
-  • MMM2 scenario with diverse units containing 1 medivac (medic), 2 marauders, and 7 marines fighting team with 1 more marauder and 1 more marine. Level: super-hard.
-  • 6h_vs_8z with 6 shooting hydralisks against 8 malee zealots. Level: super-hard.
-  • 3s5z_vs_3s6z scenario with 3 stalkers and 5 zealots against 3 stalkers and 6 zealots. Level: super-hard.
-To verify the versatility of the MACTAS algorithm, we run experiments on a diverse set of hard and super-hard scenarios.
-Every map provides the opportunity to find a unique winning strategy.
+For our experiments, we selected 9 SMACv2 maps (3 maps with 5 against 5 agents, 3 maps with 10 against 11 agents, and 3 maps with 20 against 23 agents) and 6 maps from SMAC scenarios, which are widely benchmarked in the MARL community (2c_vs_64zg, 5m_vs_6m, 27m_vs_30m, MMM2, 6h_vs_8z, 3s5z_vs_3s6z), the first 2 being hard and the latter 4 super-hard.
+On SMAC, every map provides the opportunity to find a unique winning strategy.
 The agent requires learning basic skills, such as navigating the battlefield, managing health and shield, or focusing fire.
@@
 (2019).
-In addition to classical SMAC, we benchmark the algorithms in its newer iteration, SMACv2 Ellis et al.
-(2023), which introduces new maps focused on scenario randomization.
-Each of the new maps defines the number of units in each team and the race (the same for both teams), while the exact types and starting positions are randomized according to the configuration.
+However, the standard version of SMAC is easily solved by quite simple algorithms.
+Thus, we borrow the setup from MAIC Yuan et al.
+(2022), in which the episode limits for maps have been shortened, making the benchmark challenging for newer algorithms.
+SMACv2 Ellis et al.
+(2023) is a newer iteration of SMAC, which introduces new maps focused on scenario randomization.
+Each of the new maps defines the race (the same for both teams) and the number of units in each team, while the exact types and starting positions are randomized according to the configuration.
 We use the default configuration for each scenario recommended in the SMACv2 repository.
-Our experiments focus on more challenging configurations, where the player team is outnumbered by the enemy team controlled by in-game AI, with 6 tested scenarios: zerg_5_vs_5, zerg_10_vs_11, zerg_20_vs_23, protoss_5_vs_5, protoss_10_vs_11, and protoss_20_vs_23.
-4.2. Experimental setting
+4.2 Level-based Foraging and Hallway
+We also compare MACTAS with baselines on two toy problems: Level-based Foraging (LBF) and Hallway.
+The first benchmark is a grid-world (we use $10\times 10$) with agents and food.
+Each agent and food has a level, which is randomly assigned (in our experiments, the agent’s level is at most 3 and the food’s level is at most 2).
+An agent can move in four directions, perform ”none” action, and load food.
+To collect the food, a group of agents must collectively perform the load action, and the sum of their levels must be at least as great as the food’s level.
+Agents obtain a reward for eating food, and their goal is to maximize the return in the fixed horizon (we use 50 steps).
+The maximal return is normalized to 1.
+Moreover, each agent can observe only the fixed radius around them (we use radius 2).
+Hallway is a sparse-reward coordination task with $k$ agents.
+Agent $i$ has a corridor of length $m_{i}$, is initialized in this corridor, and can move right (apart from the right end of the corridor) and left.
+There is a special common state $g$ to which each agent can move, taking a step to the left from the left end of their corridor.
+The agents’ goal is to move simultaneously to the state $g$.
+We use $k=3,m_{1}=2,m_{2}=6,m_{3}=10$.
+4.3 Experimental setting
 We borrow the basic settings from NA2Q Liu et al.
-(2023b) and extend their code with the implementation of MACTAS and MAIC.
+(2023b) and extend their code with implementations of MACTAS, MAIC, and MASIA.
 The algorithms are trained using the Double Q-Learning Hasselt et al.
 (2016) algorithm.
-Each experiment is run for a fixed number of training episode steps – 2.5 million for 2c_vs_64zg, 5m_vs_6m, MMM2, 27m_vs_30m environments and 7.5 million for 6h_vs_8z, 3s5z_vs_3s6z and all SMACv2 environments, interleaved with test episodes every 2000 steps.
+Detailed hyperparameters are available in the Appendix.
+Each experiment is run for a fixed number of training episode steps – 2.0 million for LBF and Hallway, 2.5 million for 2c_vs_64zg, 5m_vs_6m, MMM2, 27m_vs_30m SMAC environments and 7.5 million for 6h_vs_8z, 3s5z_vs_3s6z and all SMACv2 environments, interleaved with test episodes every 10000 steps.
 Each experiment is repeated for seeds $1,\dots,5$.
-We report the average percentage of wins in test games, along with the standard deviations in the Appendix.
-All hyperparameters not strictly related to the communication algorithm are shared between maps, algorithms, and mixers and are presented in Table 1.
-For MAIC communication, we employ the same hyperparameters as those used in the original paper.
-MACTAS hyperparameters
+We report the average percentage of wins in test games, along with the standard deviations, in Table 2, learning curves for the zerg race in Figure 3, and the remaining learning curves in the Appendix.
+For MAIC and MASIA communication, we employ the same hyperparameters as those used in the original papers.
+MACTAS hyperparameters.
 We notice that MACTAS is sensitive to communication hyperparameters depending on the scenario setup.
 The parameters are the number of stacked Transformer encoders and the dimension of their positional feedforward network block.
-We theorize that the depth and width of the communication channel create a trade-off between the ability to learn a complex joint policy (eg.
-kiting, hit-and-run), and the time it takes to learn an effective policy.
-We perform a hyperparameter search for each scenario, for which detailed results are included in the Appendix, and use the values presented in Table 2.
-Mixer compatibility
-Both MACTAS and MAIC are agnostic to the value function decomposition algorithm, so we compare both algorithms in combination with commonly used mixers: VDN, QMIX, and QPLEX.
-Each combination named MACTAS+VDN, MAIC+QPLEX, etc., is tested across all maps and compared separately to provide a fair comparison and isolate the effect of the mixer algorithm.
-4.3. Results
-SMAC
-The basic results are presented in Fig.
-3.
-We compare MACTAS with MAIC on 6 maps and 3 mixers, which makes 18 combinations.
-If we consider the area under the learning curve, MACTAS performs better than MAIC in 9 cases, MAIC performs better in 3 cases and there are 6 draws.
-If we consider the final performance, MACTAS yields better results 6 times, MAIC yields better results 3 times, and there are 9 draws.
-On the 2c_vs_64zg map, all algorithms except MAIC+VDN achieve the same final performance.
-In terms of the area under the learning curve MACTAS performs similarly to MAIC for the QMIX and VDN mixers and better for the QPLEX mixer.
-MACTAS+QMIX and MAIC+QMIX give jointly the best results on this map.
-On the 5m_vs_6m map MACTAS+QPLEX achieves the same final result as MAIC+QPLEX but with a better area under the learning curve.
-MACTAS+QMIX gives the same final result as MAIC+QMIX, but has a worse area under the learning curve (MAIC+QMIX drops around two million steps).
-MACTAS+VDN achieves a better final result than MAIC+VDN with a similar area under the learning curve (MAIC+VDN starts to learn faster but is unable to go beyond $60\%$ of the mean of the won battles).
-On this map MACTAS+QPLEX gives the best result.
-We observe that on the 27m_vs_30m map, our algorithm achieves a better area under the learning curve for all three mixers, and for QPLEX, it also achieves a better final performance.
-For the other mixers, the final performance is similar.
-We hypothesize that the good performance of our algorithm is because the number of parameters in our algorithm is independent of the number of agents.
-Thus, our method scales well in environments with a large number of agents.
-Overall, MACTAS+QMIX gives the best result on this map.
-On the MMM2 map, MAIC performs better for all three mixers in terms of the final result.
-However, the areas under the learning curves are similar and for MACTAS+QMIX and for MAIC+QMIX, as our algorithm starts to learn faster.
-MAIC+QPLEX performs the best on this map.
-This is the only map on which a version of our algorithm does not give the best result.
-On the 6h_vs_8z map MACTAS performs better than MAIC for the QMIX mixer in terms of the area under the learning curve and the final result.
-MACTAS+VDN gives a similar final result as MAIC+VDN, but with a better area under the learning curve.
-MACTAS+QPLEX achieves a similar final result as MAIC+QPLEX, but with a worse area under the learning curve.
-Overall, MACTAS+QMIX and MAIC+QPLEX give jointly the best result on this map.
-On the 3s5z_vs_3s6z map, MACTAS+QMIX and MACTAS+VDN give better results in terms of both the area under the learning curve and the final result than MAIC+QMIX and MAIC+VDN, respectively.
-MACTAS+QPLEX gives a similar final result to MAIC+QPLEX, but with a worse area under the learning curve.
-MACTAS+VDN performs the best on this map.
-SMACv2
-The basic results are presented in Fig.
-4.
-We compare MACTAS with MAIC and the bare mixer on 6 maps and the QMIX mixer.
-In all scenarios, our algorithm outperforms MAIC and the bare mixer.
-We observe that the differences between MACTAS+QMIX and MAIC+QMIX are similar to the differences between MAIC+QMIX and the bare mixer.
-4.4. Ablations
-There are two objectives of the ablation study.
-Firstly, we notice that the Transformer module has residual connections inside it.
+We theorize that the depth and width of the communication channel create a trade-off between the ability to learn a complex joint policy (e.g., kiting, hit-and-run) and the time it takes to learn an effective policy.
+We perform a hyperparameter search, for which detailed results are included in the Appendix, and use the values presented in Table 1.
+We use $d_{model}=64$ in the Transformer architecture (equal to the RNN size).
+Mixer compatibility.
+MACTAS, MAIC, and MASIA are agnostic to the value function decomposition algorithm, so we compare all algorithms in combination with commonly used mixers: VDN, QMIX, and QPLEX.
+Each combination, named MACTAS+VDN, MAIC+QPLEX, etc., is tested across all SMACv2 maps and compared separately to provide a fair comparison and isolate the effect of the mixer on the achieved results.
+We conduct additional experiments on SMAC, LBF, and Hallway with the QMIX mixer, as it is the default one for MAIC and MASIA.
+4.4 Results
+The basic results are presented in Table 2.
+[…]
+zerg_5_vs_5
+$\hookrightarrow$ QMIX 50.0±6.6 29.4±5.2 40.6±11.5 45.0±11.8
+$\hookrightarrow$ VDN 57.5±11.8 35.0±7.1 33.8±9.2 33.8±4.6
+$\hookrightarrow$ QPLEX 56.3±7.7 36.9±12.2 47.5±7.1 48.1±11.2
+zerg_10_vs_11
+$\hookrightarrow$ QMIX 34.4±9.9 18.8±4.9 26.3±10.5 25.6±11.8
+$\hookrightarrow$ VDN 31.3±8.3 25.0±13.4 16.3±5.1 13.1±6.8
+$\hookrightarrow$ QPLEX 36.3±10.5 1.3±1.7 19.4±18.0 19.4±2.6
+zerg_20_vs_23
+$\hookrightarrow$ QMIX 16.9±6.5 16.9±4.7 10.0±4.1 10.0±3.4
+$\hookrightarrow$ VDN 13.8±3.6 8.8±6.0 8.8±6.0 9.4±4.9
+$\hookrightarrow$ QPLEX 21.9±4.4 0.0±0.0 13.8±13.0 16.3±5.6
+terran_5_vs_5
+$\hookrightarrow$ QMIX 71.9±3.1 0.0±0.0 66.9±7.2 66.9±2.8
+$\hookrightarrow$ VDN 61.3±14.1 0.0±0.0 70.0±13.5 58.1±9.5
+$\hookrightarrow$ QPLEX 67.5±18.2 0.0±0.0 80.0±6.5 68.1±12.0
+terran_10_vs_11
+$\hookrightarrow$ QMIX 39.4±11.2 40.0±13.0 41.3±9.7 36.9±5.1
+$\hookrightarrow$ VDN 43.8±4.9 20.6±9.8 32.5±8.1 31.3±3.1
+$\hookrightarrow$ QPLEX 50.6±11.4 9.4±5.8 16.3±22.3 45.6±9.8
+terran_20_vs_23
+$\hookrightarrow$ QMIX 12.5±8.0 15.0±6.8 10.0±4.1 8.8±9.5
+$\hookrightarrow$ VDN 8.1±4.7 7.5±6.8 4.4±5.2 2.5±1.4
+$\hookrightarrow$ QPLEX 16.3±5.1 0.0±0.0 8.1±11.6 11.9±10.5
+protoss_5_vs_5
+$\hookrightarrow$ QMIX 70.6±5.2 58.8±11.8 61.9±9.2 61.3±7.5
+$\hookrightarrow$ VDN 73.8±15.2 40.6±12.3 68.1±7.1 61.3±13.2
+$\hookrightarrow$ QPLEX 70.0±7.8 56.9±13.1 70.0±6.5 60.6±11.4
+protoss_10_vs_11
+$\hookrightarrow$ QMIX 42.5±11.2 25.6±7.8 33.8±14.9 27.5±14.0
+$\hookrightarrow$ VDN 37.5±11.3 18.1±8.7 24.4±3.4 25.6±5.1
+$\hookrightarrow$ QPLEX 45.0±11.4 21.3±8.4 40.0±8.4 26.9±9.3
+protoss_20_vs_23
+$\hookrightarrow$ QMIX 15.0±10.7 13.1±6.0 15.0±3.4 10.6±6.8
+$\hookrightarrow$ VDN 9.4±9.6 6.3±2.2 5.0±2.8 2.5±1.4
+$\hookrightarrow$ QPLEX 16.9±7.8 3.8±8.4 11.9±10.2 5.6±3.4
+2c_vs_64zg
+$\hookrightarrow$ QMIX 99.4±1.4 17.5±28.8 98.1±2.8 98.8±1.7
+5m_vs_6m
+$\hookrightarrow$ QMIX 69.4±11.6 78.1±7.0 76.9±11.6 65.0±14.9
+27m_vs_30m
+$\hookrightarrow$ QMIX 78.8±12.6 93.1±8.7 87.5±7.0 75.6±11.8
+MMM2
+$\hookrightarrow$ QMIX 77.5±10.5 90.0±1.4 93.8±3.1 69.4±39.1
+6h_vs_8z
+$\hookrightarrow$ QMIX 86.9±6.8 0.0±0.0 56.9±20.1 20.0±30.4
+3s5z_vs_3s6z
+$\hookrightarrow$ QMIX 88.8±4.2 13.8±27.3 71.9±9.6 51.9±23.7
+hallway_3a_1g
+$\hookrightarrow$ QMIX 100.0±0.0 100.0±0.0 0.0±0.0 1.8±2.1
+lbf-4-2
+$\hookrightarrow$ QMIX 93.8±2.6 91.6±4.3 95.7±1.6 91.8±2.4
+Table 2: Final scores (in percent) across maps from SMACv2, SMAC, Hallway (win rate) & LBF (mean reward), for different combinations of mixers and communication algorithms.
+Bold highlights the statistically important best of the row.
+Colors are interpolated between white and green, based on the mean scaled between 0.0 and the max of each row.
+indicates NaN errors for some runs, see Appendix A.2
+SMACv2.
+We compare MACTAS with MAIC, MASIA, and the bare mixer on 9 maps and 3 different mixers.
+Our method achieves the best results on most maps.
+Moreover, MASIA fails to achieve a positive score for some scenarios, while MACTAS does not degenerate in such a manner.
+MAIC+QPLEX results in NaN errors in some runs.
+SMAC.
+We compare MACTAS with MAIC and MASIA on 6 maps, 2 hard and 4 super-hard.
+Two of the super-hard maps (6h_vs_8z, 3s5z_vs_3s6z) are particularly challenging and thus require more training steps and a longer epsilon annealing period.
+MACTAS excels on these two maps; however, it is not the best method on three simpler maps.
+Moreover, we observe that on these simpler maps, a smaller Transformer architecture (fewer layers and a smaller feedforward network) is usually required to obtain good results.
+This shows a trade-off between learning complex strategies for the most challenging scenarios and rapid learning of simpler strategies for simpler tasks.
+We also observe that MASIA fails to solve the simplest 2c_vs_64zg task, on which other methods achieve above $95\%$ of the win rate.
+LBF and Hallway.
+We observe that all methods solve the LBF task (above 90% test mean reward).
+On the Hallway benchmark, MAIC and the bare mixer fail (below the 5% test battle win rate), while MACTAS and MASIA achieve the 100% test battle win rate.
+4.5 Ablations
+We observe that the performance of MACTAS is not as outstanding on the SMAC benchmark as on the SMACv2.
+We hypothesize that the randomization introduced in SMACv2 requires finding more diverse policies and thus allows for showing the strength of MACTAS.
+To verify this hypothesis, we conduct an experiment in which, for a given seed, we fix the SMACv2 scenario and train on it MACTAS and MASIA.
+We observe that when the randomization is turned off in such a manner, MASIA can perform on a par with MACTAS (after 2.5 million training steps on protoss_5_vs_5 MACTAS achieves $96.3\pm 6.8\%$ of the battle win rate and MASIA achieves $90.0\pm 11.1\%$, on zerg_5_vs_5 MACTAS achieves $96.3\pm 4.1\%$ and MASIA achieves $90.6\pm 10.4\%$, training curves are available in the Appendix).
+We notice that the Transformer module includes residual connections.
 The question is whether the residual connection around it (the arrow between GRU and “++” in Fig.
 2) is helpful indeed.
-Figure 6 shows the impact of the residual connection on MACTAS+QMIX.
-We observe that the residual connection is crucial for the 3s5z_vs_3s6z and 6h_vs_8z maps.
-Secondly, we verify the role of our proposed exploration scheme.
-Figure 5 shows the impact of our exploration scheme for MACTAS+QMIX.
-We observe that our exploration method improves the results on the most challenging SMAC maps, which require $7.5$ million training steps: 3s5z_vs_3s6z, and 6h_8z.
-It is also indifferent on the maps: 2c_vs_64zg, 27m_vs_30m, and MMM2.
-Our exploration scheme slows down the training for the 5m_vs_6m map.
-We used softmax temperature $0.33$ for the map 3s5z_vs_3s6z and $0.25$ for the maps: 2c_vs_64zg, 5m_vs_6m, 27m_vs_30m, MMM2, and 6h_8z.
-4.5. Discussion
-[…]
+Thus, we compare the version of MACTAS without this connection to the default one.
+We observe that the absence of this residual connection hampers performance on more complex scenarios (final performance on 3s5z_vs_3s6z without this residual connection is $59.4\pm 33.2\%$, and with the connection is $88.8\pm 4.2\%$; an extended comparison is available in the Appendix).
+We compare different numbers of stacked Transformer layers for SMACv2 and conduct a broad hyperparameter search for the number of layers and the feedforward network size for SMAC in the Appendix (as we observe that MACTAS requires a dedicated architecture choice in these tasks).
+We also test several feedforward network sizes for LBF and Hallway in the Appendix.
+We observe that the architecture choice is important for the SMAC, LBF, and Hallway benchmarks, but the optimal architecture is not unique.
+Moreover, in almost all the studied cases, the versions of MACTAS outperform the bare mixer (i.e., communication turned off).
+4.6 Robustness study
+Previous experiments were performed in perfect connectivity conditions.
+In a distributed implementation, where each agent runs on separate hardware, communication must be performed over networks with potential disruptions and connectivity loss.
+We validate MACTAS in these conditions by simulating network volatility – the connectivity state $c_{i,j}$ between agents $i$ and $j$ at time $t$ is updated according to:
+$c_{i,j}^{(t)}=\begin{cases}X\sim\text{Bernoulli}(p)&\text{with probability }q\\ c_{i,j}^{(t-1)}&\text{with probability }1-q\end{cases}$ (1)
+where $q$ is the state-change probability (we use 10%) and $p$ is the probability of a functional connection (% disruption).
+Figure 4: Connectivity disruption simulation results at different $p$ values.
+Figure 4 presents the results of the study.
+Communication via the attention matrix is masked according to the above equation.
+Agents are trained according to the standard procedure, but are evaluated with a communication disturbance.
+No other changes are made to the training process or state definition; i.e., the masking matrix is not stored in the state and is not provided as an input to the model during inference or training.
+Weight updates remain unaware of any communication disruptions, making the learning conditions highly adversarial.
+MACTAS proofs immune to connectivity disruption up to $p=60\%$, while maintaining relatively strong performance up to $p=80\%$, showing a noticeable decline towards $p=100\%$ where self-attention communication starts to disturb the learning process due to the distribution shift between training and execution.
+Further reference can be found in Appendix A.3.
+5 Discussion and conclusions
+In this paper, we introduce MACTAS – a novel communication method for multi-agent reinforcement learning.
+We leverage the Transformer encoder architecture to ensure state-of-the-art performance and robust communication in heavy adversarial conditions.
+Moreover, it requires a fixed number of parameters independent of the number of agents, making it scalable to large systems.
+It can be easily switched in any multi-agent action-value function architecture.
+We evaluated our proposed module on SMACv2 and SMAC, arguably the most challenging MARL benchmarks, as well as two toy problems: LBF and Hallway.
+The experimental results show that MACTAS outperforms the baselines across most scenarios with all the tested mixers and demonstrates robustness across all tested environments.
 The modular nature of MACTAS is its great advantage.
-There are widespread variants of the Transformer code optimized for different machines.
-This makes the implementation and deployment of MACTAS in a given multi-agent system straightforward.
-Limitations
-A moderate sensitivity of the performance of our proposed communication module to its shape requires searching for its best setting in additional runs.
+There are widespread variants of the Transformer code optimized for different machines, such as PyTorch Paszke et al.
+(2019).
+This makes the implementation and deployment of MACTAS in a given MARL system straightforward.
+Limitations.
+The performance of MACTAS on SMAC, LBF, and Hallway is moderately sensitive to the shape of the communication module defined by the number of layers and the size of the hidden layer of the MLP contained there, requiring searching for its best setting in additional runs.
 This slightly slows down the deployment of MACTAS.
-5. Conclusions and Future Work
-[…]
-We observe that $\epsilon$-greedy exploration applied independently by agents in MARL leaves room for improvement since it condones the agents’ policies falling into suboptimal Nash equilibria.
-We addressed this issue by proposing a combination of $\epsilon$-greedy and Boltzmann exploration, which increased the performance of MACTAS+QMIX on the 6h_vs_8z, and the 3s5z_vs_3s6z SMAC maps.
-However, we believe that this issue requires further research and intend to undertake it in the future.
-{acks}
+Ethical Statement
+There are no ethical issues.
+Acknowledgments
 We gratefully acknowledge the Polish high-performance computing infrastructure PLGrid (HPC Center: ACK Cyfronet AGH) for providing computer facilities and support within the computational grant no.
@@
 We thank the University of Warsaw for providing access to the computing infrastructure.
-Appendix A StarCraft Multi-Agent Challenge (SMAC & SMACv2)
+Appendix A Detailed Results
+A.1 Main results
+Figures 5, 6, 7, 8, and 9 contain training curves of our experiments described in Section 4.4.
+The curves on the plots represent the mean percentage test game win rate during the algorithm training for different environments and various mixers.
+The shaded area around the curves is the standard deviation of the results among 5 different seeds (1 to 5) of the random number generator.
+For each seed, the mean percentage test win rate was calculated based on 32 independent test episodes.
+SMACv2.
+MACTAS performs the best on most maps.
+MASIA performs really good on a few maps (like protoss_20_vs_23 with the QMIX mixer), but does not outperform the bare mixer on a number of maps (like terran_5_vs_5 and protoss_5_vs_5 with all mixers, terran_10_vs_11 with QPLEX and VDN).
+MAIC performs on par with MACTAS or MASIA on a few maps, but does not visibly outperform other methods on any map.
+Moreover, MAIC+QPLEX in some runs fails to complete the training – the runs output ”NaN” in the network weights (we use the official implementation provided by the MAIC authors).
+For such cases, we report the last value before the algorithm’s failure as its output from that timestep to the end of the training.
+Figure 5: Means and standard deviations of the percentage wins in the test games during training for QMIX with MACTAS, MAIC, MASIA, and the bare mixer for the SMACv2 benchmark.
+Figure 6: Means and standard deviations of the percentage wins in the test games during training for QPLEX with MACTAS, MAIC, MASIA, and the bare mixer for the SMACv2 benchmark.
+Figure 7: Means and standard deviations of the percentage wins in the test games during training for VDN with MACTAS, MAIC, MASIA, and the bare mixer for the SMACv2 benchmark.
+SMAC.
+MACTAS with the QMIX mixer outperforms baselines on the most challenging environments 6h_vs_8z and 3s5z_vs_3s6z by a large margin.
+MASIA excels on the 27m_vs_30m and MMM2 map, but fails to solve not only the two most challenging maps, but also a simple 2c_vs_64zg.
+MAIC does not degenerate in a similar way to MASIA, but visibly outperforms MACTAS only on the MMM2 map.
+Figure 8: Means and standard deviations of the percentage wins in the test games during training for QMIX with MACTAS, MAIC, MASIA, and the bare mixer for the SMAC benchmark.
+LBF and Hallway.
+On the LBF and Hallway benchmarks with the QMIX mixer, MASIA learns the fastest, but MACTAS solves both tasks as well.
+MAIC and the bare mixer solve the LBF task, but fail to solve the Hallway benchmark.
+Figure 9: Means and standard deviations of the percentage wins in the test games during training for QMIX with MACTAS, MAIC, MASIA, and the bare mixer for the Hallway and LBF benchmarks.
+A.2 MAIC+QPLEX errors
+We observe that MAIC+QPLEX results in NaN errors in some runs.
+As mentioned before, in such cases, we report the last value before the algorithm’s failure as its output from that timestep to the end of the training.
+In Table 3, we report which seeds returned this error, after how many training steps the algorithm returned it, and the last value outputted by the algorithm.
+Map name Seed Last score Last step (million) Total environ- ment steps (million)
+terran_10_vs_11 4 0.0% 6.58 7.50
+terran_20_vs_23 2 0.0% 5.66 7.50
+terran_20_vs_23 3 0.0% 2.55 7.50
+zerg_10_vs_11 3 0.0% 6.29 7.50
+zerg_20_vs_23 1 0.0% 4.45 7.50
+Table 3: The seeds that returned the NaN error for MAIC+QPLEX on the SMACv2 benchmark.
+A.3 Robustness study
+As RL algorithms are studied in synthetic environments, many of the challenges encountered when applying these algorithms in the real world are not properly represented.
+In terms of MARL, algorithms are usually designed to maximize the speed of learning in a range of predefined environments, executing both the environment and the algorithm code together on a single machine.
+This setup can be hard to replicate in many real-world scenarios, where each agent is physically autonomous and performs all computation on embedded hardware, while communication with its peers needs to be performed over radio communication with volatility and disruptions.
+We consider robustness an important aspect of algorithm design, thus testing MACTAS against such adversary conditions.
+The exact artifacts that may arise from communication channel instability vary depending on the media and protocol, but assuming proper error detection, they all result in the agents’ inability to exchange messages for prolonged periods of time.
+To simulate this, we perturb the Transformer source mask, resulting in blocking agent pairs from exchanging information in a sequence of time steps.
+The exact algorithm is described in the main part of the paper.
+We specifically do not introduce any modifications to MACTAS learning algorithm, making it unaware of the disruptions.
+The attention masking matrix is not memorized as part of the state, and is not given to the model as an input during inference or during training.
+Figure 10 presents the results of various disruption probabilities $p$ across four SMACv1 environments.
+In one of them, with only 2 agents, 2v_vs_64zg, the disruption doesn’t impact the scores, indicating that communication is not required to solve the environment.
+In three other, MACTAS performs consistently, demonstrating strong performance for up to 60% of disruption.
+The results indicate that without any changes to the algorithm, MACTAS could be deployed in conditions of up to 60% variable time, pair-wise connection loss, and show strong learning performance.
+Figure 10: Means and standard deviations of the percentage wins for the connectivity disruption simulation at different $p$ values in the test games during training for QMIX with MACTAS for the SMAC benchmark.
+A.4 Ablations
+Residual connection.
+We study the importance of the extra residual connection in the architecture of MACTAS.
+We show the comparison of performance with and without this residual connection with QMIX as the mixer on SMAC in Figure 11.
+We observe that, on the most challenging maps (6h_vs_8z and 3s5z_vs_3s6z), the architecture with the residual connection becomes important, whereas the architecture without the residual connection remains robust on four simpler tasks.
+Figure 11: Means and standard deviations of the percentage wins in the test games during training for QMIX with MACTAS with and without the extra residual connection for the SMAC benchmark.
+Domain randomization.
+We observe that MACTAS advantage over other algorithms is most visible in difficult SMACv2 benchmarks.
+What distinguishes it from SMAC is that the positions and types of both teammates and enemies are re-randomized in every episode.
+This makes both teammate modeling (MAIC) and joint-state modeling (MASIA) harder due to a vastly larger state domain and increased variance of the inputs.
+To verify this, we set up an experiment in which we modify the SMACv2 environment behaviour so that units’ positions/types remain randomized for each global seed, but remain static for all episodes of that seed.
+This essentially removes the stochasticity of the SMACv2 environment, making the agent’s configuration static throughout the learning process.
+We observe that in this setup, MASIA can level the performance of MACTAS on the SMACv2 maps with 5 agents, on which MACTAS outperforms MASIA by a large margin when randomization is present.
+The training curves are presented in Figure 12.
+We conclude that the advantages of MACTAS are particularly evident in complex tasks, where the agent must adapt to different scenarios rather than learn a policy for a fixed scenario.
+Figure 12: Means and standard deviations of the percentage wins in the test games during training for QMIX with MACTAS and MASIA with full randomization and fixed episode generation for two SMACv2 maps.
+A.5 QPLEX variants
+We note that the QPLEX mixer Wang et al.
+[2021] is presented in two variants: with the MLP and the attention architecture.
+For MACTAS, we chose the MLP variant and compared it with the MLP variant of QPLEX.
+However, MASIA and MAIC use the attention variant, and we followed their choice in the main experiments.
+We study how MACTAS+QPLEX and QPLEX perform if we change the architecture to the attention one.
+The results are presented in Figure 13.
+We observe that for MACTAS the attention variant performs similarly on the most maps, better on the protoss_20_vs_23 and worse on the zerg_20_vs_23 and terran_20_vs_23 maps.
+The attention variant for the QPLEX usually performs not worse than the MLP variant, but fails on the terran_20_vs_23 map.
+We observe that MACTAS+QPLEX with the attention architecture outperforms QPLEX with the attention architecture, MASIA, and MAIC.
+Figure 13: Means and standard deviations of the percentage wins in the test games during training for MACTAS+QPLEX and QPLEX with the MLP and the attention QPLEX architecture for the SMACv2 maps.
+Appendix B Sensitivity analysis
+We observe that MACTAS is moderately sensitive to the architecture choice of the Transformer encoder, namely, the number of stacked layers of the Transformer encoder and the dimension of the feedforward network (we use in all experiments $d_{model}=64$, as it should be equal to the RNN hidden dimension).
+We use a consistent architectural choice for the most challenging benchmarks (all SMACv2 maps and two SMAC maps: 3s5z_vs_3s6z, 6h_vs_8z), and the toy problems (LBF and Hallway).
+For the remaining four SMAC maps, we use the same architecture for 27m_vs_30m, 5m_vs_6m, and a separate one 2c_vs_64zg and MMM2.
+We present a study of the robustness of MACTAS across different architectural choices with QMIX as the mixer.
+We denote by ”lxx dimffyy” the architecture with $x$ layers and the size of the feedforward network hidden layer equal to $y$.
+B.1 SMACv2
+We show the sensitivity of MACTAS to the number of Transformer encoder layers with the size of the feedforward hidden layer equal to 512 in Figure 14.
+On the SMACv2 benchmark, we observe that MACTAS performs well with the default architecture.
+On some maps (terran_10_vs_11, protoss_20_vs_23), one can slightly improve the results by taking a different number of layers.
+All tested number of layers perform not worse than the bare mixer, apart from the case of one layer on the zerg_20_vs_23 map.
+Figure 14: MACTAS; sensitivity of the results to architecture size with the QMIX mixer on the SMACv2 benchmark.
+B.2 SMAC
+As MACTAS is more sensitive to the architectural choice on SMAC, we show the whole hyperparameter search.
+We test the one, two, three, and four stacked Transformer encoder layers and the dimension of the Transformer feedforward network equal to 128, 256, 512, and 1024.
+The results are presented in Figures 15, 16, 17, and 18.
+We observe that MACTAS performs not worse than the bare mixer in almost all tested configurations and scenarios.
+On the 2c_vs_64zg and 5m_vs_6m maps, most of the setups perform equally well.
+On the remaining maps, the architectural choice is important, but there are alternative settings performing similarly to the optimal architecture (for example, l1dimff1024, l3dimff1024, l4dimff1024 on the 6h_vs_8z map; l2dimff512 on the 3s5z_vs_3s6z map).
+Figure 15: MACTAS; sensitivity of the results to architecture size with the QMIX mixer and the size of the feedforward connection equal to 128 and optimal on the SMAC benchmark.
+Figure 16: MACTAS; sensitivity of the results to architecture size with the QMIX mixer and the size of the feedforward connection equal to 256 and the optimal architecture on the SMAC benchmark .
+Figure 17: MACTAS; sensitivity of the results to architecture size with the QMIX mixer and the size of the feedforward connection equal to 512 and the optimal architecture on the SMAC benchmark .
+Figure 18: MACTAS; sensitivity of the results to architecture size with the QMIX mixer and the size of the feedforward connection equal to 1024 and the optimal architecture on the SMAC benchmark.
+B.3 LBF and Hallway
+On the simple tasks, MACTAS requires a small Transformer architecture to learn fast.
+Thus, for the toy problems, we choose the one-layer architecture.
+We test the Transformer encoder with feedforward network widths of 32, 64, and 128.
+The results are presented in Figure 9.
+We observe that one can obtain even better results on the Hallway by selecting a slightly different architecture, but we aim to keep the architecture consistent across the toy problems.
+Figure 19: MACTAS; sensitivity of the results to architecture size with the QMIX mixer and one-layer architecture on the Hallway and LBF benchmarks.
+Appendix C StarCraft Multi-Agent Challenge (SMAC & SMACv2)
 As a benchmark for MACTAS, we chose the StarCraft Multi-Agent Challenge (SMAC) Samvelyan et al.
-(2019) and its newer version SMACv2 Ellis et al.
-(2023), which are simplified versions of the StarCraft II game.
+[2019] and its newer version SMACv2 Ellis et al.
+[2023], which are simplified versions of the StarCraft II game.
 The framework exists in two versions – SMAC (code available at https://github.com/oxwhirl/smac), which introduced predefined deterministic scenarios that often include purposeful selection of units or obstacles placement, and SMACv2 (code available at https://github.com/oxwhirl/smacv2), which added new randomized scenarios based on plain maps and randomized units generation.
@@
 The process for running SMAC is described in our repository.
-Appendix B Detailed Results
-Figures 7, 8, and 9 contain more detailed results of our experiments described in Section 4.3.
-The curves on the plots represent the mean percentage test game win rate during the algorithm training for 6 different environments and various mixers.
-The shaded area around the curves is the standard deviation of the results among 5 different seeds (1 to 5) of the random number generator.
-For each seed, the mean percentage test win rate was calculated based on 32 independent test episodes.
-We observe robustness of MACTAS with the QMIX mixer – on the most challenging environments 6h_vs_8z and 3s5z_vs_3s6z, it outperforms MAIC+QMIX by a large margin; on 27m_vs_30m, it learns much faster; on 2c_vs_64zg and 3s5z_vs_3s6z, it performs similarly, and only slightly worse on the MMM2 map.
-For the QPLEX mixer, our method performs much better than MAIC on the 2c_vs_64zg and 27m_vs_30m maps, slightly better on the 5m_vs_6m, slightly worse on MMM2, learns slower but achieves similar final result on the 6h_vs_8z and 3s5z_vs_3s6z maps.
-Our method combined with the VDN mixer gives significantly better results than MAIC on the 3s5z_vs_3s6z map, slightly better on 6h_vs_8z, 27m_vs_30m, 5m_vs_6m, marginally better on the 2c_vs_64zg map, and marginally worse on the MMM2 map.
-Figure 7. Means and standard deviations of the percentage wins in the test games during training for QMIX with MAIC, MACTAS, and bare.
-Figure 8. Means and standard deviations of the percentage wins in the test games during training for QPLEX with MAIC, MACTAS, and bare.
-Figure 9. Means and standard deviations of the percentage wins in the test games during training for VDN with MAIC, MACTAS, and bare.
-Appendix C Tactics used by MACTAS
-The SMAC challenge forces agents to learn various tactics, such as attacking from a distance with melee units (’kiting’) or using a movement advantage (’hit and run’).
-In this section, we provide an overview of the tactics used by our model based on our understanding of StarCraft II.
-We observed differences in the behavior of our model depending on the random seed.
-The description of tactics is based on replays of our models for a seed value of 1.
-C.1. 2c_vs_64zg
-Figure 10. 2c_vs_64zg replay.
-In this scenario, agents control 2 colossi units.
-Colossi can move through hills between the high and the low ground.
-Against them, 64 enemy zerglings are deployed, which are fast ground units.
-Zerglings are most effective when they surround an opposite unit.
-In the scenario 2c_vs_64zg, the colossi start on the hill, while the zerglings are split into two groups of 32 on the left and the right, both of them on the low ground.
-Figure 10 presents the following tactics: colossi attack the right group of zerglings and wait for the left group to climb the hill.
-Then, colossi use the ability to traverse the hill and attack the remaining zerglings, which are forced to run around the hill.
-In this way, the colossi avoid being surrounded and thereby defeated.
-C.2. 27m_vs_30m and 5m_vs_6m
-Figure 11. 27m_vs_30m replay.
-Figure 12. 5m_vs_6m replay.
-The 27m_vs_30m and the 5m_vs_6m scenarios consist of marine units only.
-They are similar in the way that the opposing force numerically outnumbers the agents under control.
-In Figure 11, the agent’s units place themselves in a semicircle formation, also known as flanking, which maximizes the number of shooting units, while the enemies block each other.
-Units controlled by MACTAS focus fire on enemy units with lower health to quickly reduce the number of opponents.
-Furthermore, units with low health retreat and continue fighting, while avoiding being shot.
-This behavior is presented in Figure 12: Agents with low health retreat when others focus fire on one enemy unit.
-The winning strategy is similar for both 5m_vs_6m and 27m_vs_30m.
-5m_vs_6m is labeled as a hard benchmark, while 27m_vs_30m is classified as super-hard.
-The percentage imbalance between units is higher in 5m_vs_6m.
-However, 27m_vs_30m has a larger number of units that need to cooperate.
-C.3. MMM2
-Figure 13. MMM2 replay.
-The MMM2 is a scenario with a diverse set of units containing 1 medivac (medic), 2 marauders, and 7 marines fighting against a team with 1 more marauder and 1 more marine.
-Medivac cannot attack, but can heal units as long as it has energy.
-MACTAS ground units focus fire on enemy units and eliminate them one by one, and fall back when their health is low.
-The enemy medivac does not change healed unit, while the MACTAS’ medivac changes healed unit to the one with a critically low health or to the currently attacked unit.
-Units trained with MACTAS move medivac to the front line and focus enemy fire on it, and when its health level is low, retreat it.
-MACTAS’ medivac taking enemy fire secures outnumbered marines and marauders.
-Typically, human players try to save their medivacs because they are used for healing units between fights during a regular game.
-Only under certain conditions do they sacrifice their medivacs; for example, when they are breaking through an opponent’s line of defense to focus enemy fire on the medivacs and away from damage-dealing units.
-In this scenario, we can observe similar behavior in the MACTAS medivac control, which results in winning combat against an opponent who has a numerical advantage.
-C.4. 6h_vs_8z
-Figure 14. 6h_vs_8z replay.
-Agents control 6 range shooting units, hydralisks, while the enemy controls 8 melee zealots.
-In Figure 14, we can see a tactic called “kiting” – hydralisks split into two groups and are attacking zealots from distances.
-Enemy zealots attack the nearest unit – our agents learned to move enough to change the focus of approaching units and then continue striking.
-Notably, the zealot units within this scenario are not upgraded with the “charge” ability, which would otherwise allow them to have a significant increase in movement speed and easily reach hydralisks.
-C.5. 3s5z_vs_3s6z
-Figure 15. 3s5z_vs_3s6z replay.
-MACTAS controls 3 stalkers and 5 zealots, while the enemy has one more zealot.
-The stalkers are fast, range-attacking units, but with weak defense.
-On the other hand, zealots have a good defense, but they need to get close to the opponent’s unit.
-The tactic in this scenario is interesting.
-Using one stalker as a decoy, MACTAS first defeats the enemy stalkers and then proceeds to eliminate the rest of the enemy force.
-One agent, a stalker, sacrifices himself and focuses a few enemy zealots on himself.
-In Figure 15, we can see 5 enemy zealots following the stalker marked as “Local Player”.
-The remaining MACTAS’ 2 stalkers and 5 zealots attack 3 enemy stalkers and 1 zealot with an advantage.
-This move results in the elimination of enemy stalkers and the zealot.
-In the end, MACTAS’ units eliminate zealots lured by the sacrificed stalker.
-Appendix D Sensitivity analysis
-Figure 16. MACTAS; sensitivity of the results to architecture size with the QMIX mixer.
-Figure 17. MACTAS; sensitivity of the results to architecture size with the QPLEX mixer.
-Figure 18. MACTAS; sensitivity of the results to architecture size with the VDN mixer.
-Figure 19. MACTAS; sensitivity of the results to the dropout parameter in the Transformer block. The default value of this parameter is $10\%$.
-Figure 20. MACTAS; sensitivity of the top 2 exploration methods to the temperature coefficient.
-We evaluated the choice of the number of Transformer encoder layers and the width of the Transformer feedforward layer for MACTAS with QMIX, QPLEX, and VDN as a mixer on six SMAC(v1) maps: 2c_vs_64zg, 5m_vs_6m, 27m_vs_30m, MMM2, 6h_vs_8z, and 3s5z_vs_3s6z.
-We experimented with the dropout of the Transformer encoder on all six maps.
-We also assessed the sensitivity of our new top $k$ exploration technique to the softmax temperature in the case of MACTAS+QMIX and $k=2$ on all six maps.
-[…]
-Figure 19 shows how the dropout parameter affects the performance of MACTAS.
-We observe that the absence of dropout harms the performance; however, the choice of the dropout fraction is not crucial when set in a reasonable interval; relevant differences can be observed only on the 3s5z_vs_3s6z map.
-Figure 20 shows the impact of the softmax temperature parameter on the proposed exploration scheme.
-We observe that this method is sensitive to this parameter.
-For the 2c_vs_64zg, 5m_vs_6m, 27m_vs_30m, MMM2, and 6h_vs_8z maps, the low temperatures ($0.25$, $0.33$) work best with better results for the $0.25$ temperature.
-We can observe an interesting situation on the 3s5z_vs_3s6z map.
-The best results are obtained for the $0.33$ temperature, with the temperature $4$ being the second-best, and the temperature $0.25$ performing quite poorly, which suggests that, in general, the relationship between the temperature value and performance may not be monotonic.
-For the most challenging maps, which require 7.5 million training steps, the top 2 exploration schemes outperform the $\epsilon$-greedy exploration, which is equivalent to choosing 0 as the temperature.
-In other cases, the new exploration scheme does not outperform the $\epsilon$-greedy exploration; however, except for the case of the 5m_vs_6m map, where training is slowed down, the method remains indifferent.
-Appendix E Software
-Table 3 contains a list of key programming libraries used during our experiments.
-The complete list of programming libraries is available in the MACTAS’ official code repository in the file requirements.txt.
+Appendix D Setup hyparparameters
+All hyperparameters not strictly related to the communication algorithm are shared between maps, algorithms, and mixers and are presented in Table 4.
+Hyperparameter Value
+Batch size 32
+Test episodes 32
+Replay buffer size 5000
+Discount factor 0.99
+Start epsilon 1.0
+Finish epsilon 0.05
+Anneal steps 50,0001 or 100,0002
+RNN units 64
+Optimizer RMSProp
+Learning rate 0.0005
+Table 4: Basic experimental settings and hyperparameters. 1 for LBF, Hallway, and 2.5-million step SMAC maps. 2 for SMACv2 and 7.5 million-step SMAC maps.
 Library name Version
@@
 torch 2.5.1
-Table 3. Essential libraries in programming environment.
-Appendix F Hardware
+Table 5: Essential libraries in programming environment.
 Cluster name GPU device GPU driver version CPU device Operating system Total RAM
@@
 Cluster 3 NVIDIA A100 PCIe 80GB 565.57.01 AMD EPYC 7713 64-Core Processor Linux-6.8.0-64-generic-x86_64-with-glibc2.39 2048 GB
-Table 4. Specification of the computational clusters used in the experiments.
+Table 6: Specification of the computational clusters used in the experiments.
+Appendix E Tactics used by MACTAS
+The SMAC challenge forces agents to learn various tactics, such as attacking from a distance with melee units (’kiting’) or using a movement advantage (’hit and run’).
+In this section, we provide an overview of the tactics used by our model based on our understanding of StarCraft II.
+We observed differences in the behavior of our model depending on the random seed.
+The description of tactics is based on replays of our models for a seed value of 1.
+We describe tactics on the following maps from SMACv1:
+  • 2c_vs_64zg with 2 colossus units leveraging terrain advantage against 64 small enemy zerglings. Level: hard.
+  • 5m_vs_6m with 5 cooperating marines and 6 enemy marines. Level: hard.
+  • 27m_vs_30m big battle of 27 marines conquering 30 marines. Level: super-hard.
+  • MMM2 scenario with diverse units containing 1 medivac (medic), 2 marauders, and 7 marines fighting team with 1 more marauder and 1 more marine. Level: super-hard.
+  • 6h_vs_8z with 6 shooting hydralisks against 8 malee zealots. Level: super-hard.
+  • 3s5z_vs_3s6z scenario with 3 stalkers and 5 zealots against 3 stalkers and 6 zealots. Level: super-hard.
+E.1 2c_vs_64zg
+Figure 20: 2c_vs_64zg replay.
+In this scenario, agents control 2 colossi units.
+Colossi can move through hills between the high and the low ground.
+Against them, 64 enemy zerglings are deployed, which are fast ground units.
+Zerglings are most effective when they surround an opposite unit.
+In the scenario 2c_vs_64zg, the colossi start on the hill, while the zerglings are split into two groups of 32 on the left and the right, both of them on the low ground.
+Figure 20 presents the following tactics: colossi attack the right group of zerglings and wait for the left group to climb the hill.
+Then, colossi use the ability to traverse the hill and attack the remaining zerglings, which are forced to run around the hill.
+In this way, the colossi avoid being surrounded and thereby defeated.
+E.2 27m_vs_30m and 5m_vs_6m
+Figure 21: 27m_vs_30m replay.
+Figure 22: 5m_vs_6m replay.
+The 27m_vs_30m and the 5m_vs_6m scenarios consist of marine units only.
+They are similar in the way that the opposing force numerically outnumbers the agents under control.
+In Figure 21, the agent’s units place themselves in a semicircle formation, also known as flanking, which maximizes the number of shooting units, while the enemies block each other.
+Units controlled by MACTAS focus fire on enemy units with lower health to quickly reduce the number of opponents.
+Furthermore, units with low health retreat and continue fighting, while avoiding being shot.
+This behavior is presented in Figure 22: Agents with low health retreat when others focus fire on one enemy unit.
+The winning strategy is similar for both 5m_vs_6m and 27m_vs_30m.
+5m_vs_6m is labeled as a hard benchmark, while 27m_vs_30m is classified as super-hard.
+The percentage imbalance between units is higher in 5m_vs_6m.
+However, 27m_vs_30m has a larger number of units that need to cooperate.
+E.3 MMM2
+Figure 23: MMM2 replay.
+The MMM2 is a scenario with a diverse set of units containing 1 medivac (medic), 2 marauders, and 7 marines fighting against a team with 1 more marauder and 1 more marine.
+Medivac cannot attack, but can heal units as long as it has energy.
+MACTAS ground units focus fire on enemy units and eliminate them one by one, and fall back when their health is low.
+The enemy medivac does not change healed unit, while the MACTAS’ medivac changes healed unit to the one with a critically low health or to the currently attacked unit.
+Units trained with MACTAS move medivac to the front line and focus enemy fire on it, and when its health level is low, retreat it.
+MACTAS’ medivac taking enemy fire secures outnumbered marines and marauders.
+Typically, human players try to save their medivacs because they are used for healing units between fights during a regular game.
+Only under certain conditions do they sacrifice their medivacs; for example, when they are breaking through an opponent’s line of defense to focus enemy fire on the medivacs and away from damage-dealing units.
+In this scenario, we can observe similar behavior in the MACTAS medivac control, which results in winning combat against an opponent who has a numerical advantage.
+E.4 6h_vs_8z
+Figure 24: 6h_vs_8z replay.
+Agents control 6 range shooting units, hydralisks, while the enemy controls 8 melee zealots.
+In Figure 24, we can see a tactic called “kiting” – hydralisks split into two groups and are attacking zealots from distances.
+Enemy zealots attack the nearest unit – our agents learned to move enough to change the focus of approaching units and then continue striking.
+Notably, the zealot units within this scenario are not upgraded with the “charge” ability, which would otherwise allow them to have a significant increase in movement speed and easily reach hydralisks.
+E.5 3s5z_vs_3s6z
+Figure 25: 3s5z_vs_3s6z replay.
+MACTAS controls 3 stalkers and 5 zealots, while the enemy has one more zealot.
+The stalkers are fast, range-attacking units, but with weak defense.
+On the other hand, zealots have a good defense, but they need to get close to the opponent’s unit.
+The tactic in this scenario is interesting.
+Using one stalker as a decoy, MACTAS first defeats the enemy stalkers and then proceeds to eliminate the rest of the enemy force.
+One agent, a stalker, sacrifices himself and focuses a few enemy zealots on himself.
+In Figure 25, we can see 5 enemy zealots following the stalker marked as “Local Player”.
+The remaining MACTAS’ 2 stalkers and 5 zealots attack 3 enemy stalkers and 1 zealot with an advantage.
+This move results in the elimination of enemy stalkers and the zealot.
+In the end, MACTAS’ units eliminate zealots lured by the sacrificed stalker.
+Appendix F Software
+Table 5 contains a list of key programming libraries used during our experiments.
+The complete list of programming libraries is available in the MACTAS’ official code repository in the file requirements.txt.
+Appendix G Hardware
 Here we present the hardware used in our experiments. We utilized the NVIDIA Ampere GPU architecture across three computational clusters. The specific setup is presented in Table 6.